### v2

# Polynomials Counting Group Colorings in Graphs

Abstract
Jaeger et al. in 1992 introduced group coloring as the dual concept to group connectivity in graphs. Let $\Gamma$ be an Abelian group, $f:E(G)\to\Gamma$ and $D$ an orientation of a graph $G$. A vertex coloring $c:V(G)\to\Gamma$ is a $(\Gamma,f)$-coloring if $c(v)-c(u)\neq f(e)$ for each edge $e=uv$, where the corresponding arc $D(e)=(u,v)$ is directed from $u$ to $v$. We introduce the concept of $\alpha$-compatible graphs and define the cycle-assigning polynomial $P(G,\alpha;k)$ at $k$ in terms of $\alpha$-compatible spanning subgraphs, where $\alpha$ is an assigning of $G$ from its cycles to $\{0,1\}$. We prove that the cycle-assigning polynomial $P(G,\alpha;k)$ equals the number of $(\Gamma,f)$-colorings for any Abelian group $\Gamma$ of order $k$ and $f:E(G)\to\Gamma$ such that the assigning $\alpha_{D,f}$ induced by $f$ equals $\alpha$. In particular, $P(G,\alpha;k)$ is the classical chromatic polynomial if $\alpha$ is the zero function. Furthermore, we introduce the concept of $\alpha$-compatible broken cycles and interpret $P(G,\alpha;k)$ in terms of $\alpha$-compatible spanning subgraphs that do not contain $\alpha$-compatible broken cycles. This implies that the absolute value of the coefficient of $k^{r(G)-i}$ in $P(G,\alpha;k)$ equals the number of $\alpha$-compatible spanning subgraphs that have $i$ edges and contain no $\alpha$-compatible broken cycles, which generalizes the Whitney’s Broken Cycle Theorem. Based on the combinatorial explanation, we establish a unified order-preserving relation from assignings to cycle-assigning polynomials. Finally, we show that for any loopless graph $G$, the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ are nonzero and alternate in sign, and further conjecture that the sequence of absolute values of its coefficients is unimodal and log-concave.
Keywords: Graph coloring, chromatic polynomial, $\alpha$-compatible graph, broken cycle
Mathematics Subject Classifications: 05C31, 05C15
1 Introduction
The graphs $G$ with vertex set $V(G)$ and edge set $E(G)$ considered in this paper are finite, allowing multiple edges and loops. A more general setting and background on graphs can be found in the book [4]. An orientation of a graph is an assignment of direction to each edge. In this context, each oriented edge is referred to as an arc. An oriented graph arises from a graph $G$ together with a particular orientation $D$, denoted by $D(G):=\big{(}V(D),A(D)\big{)}$, where $V(D)=V(G)$ and $A(D)$ is the collection of arcs with respect to the orientation $D$.
Graph coloring is a significant subfield in graph theory that originated in the middle of the 19th century with the famous Four Color Problem. A $k$-vertex coloring, or simply a $k$-coloring, of $G$ is a mapping $c:V(G)\to S$, where $S$ is a set of distinct $k$ colors, typically $S=\{1,2,...,k\}$. A vertex coloring is said to be proper if no two adjacent vertices share the same color.
We consider finite additive Abelian groups $\Gamma$ with the identity 00 as coloring sets in this paper. Jaeger, Linial, Payan and Tarsi [12] introduced the concept of group coloring as the dual concept of group connectivity of graphs in 1992. Group connectivity and group colorings of graphs are nicely surveyed in [16]. Let $f:E(G)\to\Gamma$ be a function. With an orientation $D$ of $G$, a vertex coloring $c:V(G)\to\Gamma$ is called a $(\Gamma,f)$-coloring if $c(v)-c(u)\neq f(e)$ for each edge $e=uv$, where the corresponding arc $D(e)=(u,v)$ is directed from $u$ to $v$. When $f=0$, $(\Gamma,0)$-coloring is exactly the ordinary proper coloring. Furthermore, $G$ is said to be $\Gamma$-colorable under the orientation $D$ if for any function $f:E(G)\to\Gamma$, $G$ always admits a $(\Gamma,f)$-coloring. Since then, topics related to group coloring problems have attracted a lot of research attentions, such as [14, 19, 20, 21, 15, 17, 18]. Their primary focus is on the (weak) group chromatic number and related group coloring problems, such as whether a graph that is $\Gamma_{1}$-colorable for some Abelian group $\Gamma_{1}$ is also $\Gamma_{2}$-colorable for any Abelian group $\Gamma_{2}$ of the same order. However, another fundamental problem of counting $(\Gamma,f)$-colorings has been overlooked. In this paper, we introduce the cycle-assigning polynomial that counts the number of $(\Gamma,f)$-colorings of graphs, and investigate its fundamental properties, particularly focusing on its coefficients.
A fundamental problem in graph theory is determining the chromatic number of a graph, which refers to the smallest integer $k$ such that the graph admits a proper coloring with $k$ colors. Counting proper colorings can be seen as another facet of this problem. In an attempt to address the long-standing Four Color Problem, Birkhoff [2] discovered that the number of proper colorings of a planar $G$ with $k$ colors is a polynomial in $k$, called the chromatic polynomial. Twenty years later, Whitney [26, 27] extended the concept of the chromatic polynomial to general graphs and gave a combinatorial description of its coefficients by introducing the broken cycle, now famously known as the Whitney’s Broken Cycle Theorem. Since then, various generalizations of this theorem have been developed across different mathematical structures, such as matroids [5, 8], lattices [3, 24], hyperplane arrangements [22], hypergraphs [6]. See [7] for more related work. Motivated by their work, it is natural to ask the two questions:
  • Is there a polynomial function of $|\Gamma|$ that generalizes the chromatic polynomial and counts $(\Gamma,f)$-colorings in graphs?
  • Can we provide a combinatorial interpretation for the coefficients of the polynomial function?
To answer the first question, we propose the concept of $\alpha$-compatible graphs. A cycle $C$ of $G$ is a connected 2-regular subgraph of $G$. The corresponding edge set $E(C)$ is referred to as a circuit of $G$. Let $\mathcal{C}(G)$ be the family of cycles of $G$. An assigning $\alpha$ of $G$ is a mapping from $\mathcal{C}(G)$ to the set $\{0,1\}$. In particular, we define $\alpha=0$ when $G$ does not contain cycles, i.e., $\mathcal{C}(G)=\emptyset$.
[…]
Let $\alpha$ be an assigning of $G$. A graph $G$ is said to be $\alpha$-compatible if $\alpha(C)=0$ for every cycle $C$ in $G$. Furthermore, a subgraph $H$ of $G$ is called an $\alpha$-compatible subgraph if the restriction of $\alpha$ to $H$ satisfies $\alpha(C)=0$ for all cycles $C$ in $H$. In this sense, $H$ is $\alpha_{H}$-compatible, where $\alpha_{H}$ is the restriction of $\alpha$ to $H$ defined by $\alpha_{H}(C):=\alpha(C)$ for all $C\in\mathcal{C}(H)$.
Let $f:E(G)\to\Gamma$, and $P\big{(}D(G),f;\Gamma\big{)}$ be the number of $(\Gamma,f)$-colorings of $G$ with an orientation $D$. We first provide an explicit formula for the counting function $P\big{(}D(G),f;\Gamma\big{)}$ in terms of the $\alpha$-compatible spanning subgraph expansion in Theorem 2.1, where $\alpha$ equals the admissible assigning $\alpha_{D,f}$ induced by $f$. The detailed definition of an admissible assigning can be found at the end of the second paragraph in Section 2. Inspired by this expression, we define a cycle-assigning polynomial in terms of $\alpha$-compatible spanning subgraphs. It turns out that cycle-assigning polynomials precisely count $(\Gamma,f)$-colorings of graphs, as shown in Theorem 2.2.
[…]
For any assigning $\alpha$ of $G$, the cycle-assigning polynomial of $G$ in $k$ is defined as
$P(G,\alpha;k):=\sum_{\begin{subarray}{c}\mbox{$H$ is an $\alpha$-compatible}\\ \mbox{spanning subgraph of $G$}\end{subarray}}(-1)^{|E(H)|}k^{c(H)},$
where $c(H)$ is the number of components of $H$.
To address the second question, we introduce the concept of an $\alpha$-compatible broken cycle as an extension of the usual broken cycle. In Theorem 3.1, we demonstrate that the unsigned coefficients of cycle-assigning polynomials can be expressed in terms of $\alpha$-compatible spanning subgraphs that do not contain $\alpha$-compatible broken cycles. This extends the Whitney’s celebrated Broken Cycle Theorem [26]. Based on the counting interpretation, we further establish a unified order-preserving relation from assignings to cycle-assigning polynomials when both are naturally ordered in Theorem 4.1. In addition, we also provide a new perspective on addressing certain group coloring problems in graphs. A detailed discussion will appear at the end of Section 4.
The paper is organized as follows. Section 2 presents an explicit formula for calculating $(\Gamma,f)$-colorings (see Theorem 2.1), and further shows that the number of $(\Gamma,f)$-colorings with the function $f$ corresponding to an assigning is evaluated by the cycle-assigning polynomial of $|\Gamma|$ in Theorem 3.1. In Section 3, we show that the cycle-assigning polynomial $P(G,\alpha;k)$ can be interpreted in terms of $\alpha$-compatible spanning subgraphs that do not contain $\alpha$-compatible broken cycles in Theorem 3.1, which generalizes the Whitney’s Broken Cycle Theorem and is called the Generalized Whitney’s Broken Cycle Theorem. Section 4 utilizes the Generalized Whitney’s Broken Cycle Theorem to establish a unified order-preserving relation from assignings to cycle-assigning polynomials in Theorem 4.1, demonstrates that for any loopless graph $G$, the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ are nonzero and alternate in sign Section 4, and conjecture that the absolute values of its coefficients form a unimodal and log-concave sequence in Section 4.
2 Cycle-assigning polynomials
In this section, we focus on the study of cycle-assigning polynomials, exploring their structural properties and their role in counting $(\Gamma,f)$-colorings. Let us review some necessary notations and definitions on graphs. A graph $H$ is a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. When $V(H)=V(G)$, the subgraph $H$ of $G$ is referred to as a spanning subgraph. For an edge subset $S\subseteq E(G)$, $G-S$ is the subgraph obtained by removing the edges in $S$. Denote by $G\langle S\rangle$ the spanning subgraph with edges $S$, that is, $G\langle S\rangle=G-(E(G)\smallsetminus S)$. Conversely, suppose $H$ is a proper subgraph containing no edge $e\in E(G)$, we denote by $H+e$ the subgraph of $G$ by adding the edge $e$. In addition, if $e$ is not a loop of $G$, the graph $G/e$ is obtained by contracting $e$, that is, identifying the ends of $e$ to create a single new vertex and then removing $e$.
Let $\mathcal{O}(G)$ be the set of all orientations of $G$. Unless otherwise stated, we always choose an arbitrary but fixed reference orientation $D$ for $G$. For each cycle $C$ of $G$, define a map $\sigma_{D(C)}$ from the circuit $E(C)$ to the set $\{1,-1\}$ as
$\sigma_{D(C)}(e):=\begin{cases}1,&\mbox{if the direction of the arc $D(e)$ is clockwise with respect to $C$};\\ -1,&\mbox{if the direction of the arc $D(e)$ is counter-clockwise with respectto $C$}.\end{cases}$
Notably, each function $f:E(G)\to\Gamma$ automatically induces an assigning $\alpha_{D,f}$ on $\mathcal{C}(G)$ such that for each cycle $C$ of $G$,
$\alpha_{D,f}(C)=\begin{cases}0,&\mbox{ if }\sum_{e\in E(C)}\sigma_{D(C)}(e)f(e)=0;\\ 1,&\mbox{ otherwise }.\end{cases}$
We say that $\alpha=\alpha_{D,f}$ is $D$-admissible, or simply admissible.
Recall from Section 1 that the counting function $P\big{(}D(G),f;\Gamma\big{)}$ is the number of $(\Gamma,f)$-colorings of $G$ associated with the orientation $D$ and a function $f:E(G)\to\Gamma$. To derive an explicit formula for the counting function $P\big{(}D(G),f;\Gamma\big{)}$, we need to introduce the Möbius Inversion Formula in [4, Theorem 2.25]. The Möbius Inversion Formula states that if the function $g:2^{E(G)}\to\Gamma$ is defined as
$g(S):=\sum_{S\subseteq X\subseteq E(G)}h(X)\quad\mbox{ for }\quad\forall\,S\subseteq E(G),$
then, for all $S\subseteq E(G)$,
$h(S)=\sum_{S\subseteq X\subseteq E(G)}(-1)^{|X|-|S|}g(X),$ (2.1)
where $h$ is a function from the family $2^{E(G)}$ of all edge subsets of $G$ to $\Gamma$.
The following result shows that the counting function $P\big{(}D(G),f;\Gamma\big{)}$ can be expressed in terms of the $\alpha_{D,f}$-compatible spanning subgraph expansion.
Theorem 2.1 (Counting Formula).
Let $f:E(G)\to\Gamma$ and $\alpha=\alpha_{D,f}$. Then
$P\big{(}D(G),f;\Gamma\big{)}=\sum_{\begin{subarray}{c}\mbox{$H$ is an $\alpha$-compatible}\\ \mbox{spanning subgraph of $G$}\end{subarray}}(-1)^{|E(H)|}|\Gamma|^{c(H)}.$
Proof.
For any $S\subseteq E(G)$, let $C_{S}\big{(}D(G),f;\Gamma\big{)}$ be the number of vertex colorings $c:V(G)\to\Gamma$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$, and $P_{S}\big{(}D(G),f;\Gamma\big{)}$ be the number of vertex colorings $c:V(G)\to\Gamma$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$ and $c(v)-c(u)\neq f(e)$ for each edge $e=uv\in E(G)\smallsetminus S$ with $D(e)=(u,v)$. Then we have
$C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}=C_{S}\big{(}D(G),f;\Gamma\big{)},\quad\quad P\big{(}D(G),f;\Gamma\big{)}=P_{\emptyset}\big{(}D(G),f;\Gamma\big{)}$
and
$C_{\emptyset}\big{(}D(G),f;\Gamma\big{)}=\sum_{S\subseteq E(G)}P_{S}\big{(}D(G),f;\Gamma\big{)}.$
It follows from (2.1) that
$P\big{(}D(G),f;\Gamma\big{)}=\sum_{S\subseteq E(G)}(-1)^{|S|}C_{S}\big{(}D(G),f;\Gamma\big{)}=\sum_{S\subseteq E(G)}(-1)^{|S|}C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}.$ (2.2)
In addition, [13, Proposition 2.2] demonstrates that $G$ admits a vertex coloring $c$ such that $c(v)-c(u)=f(e)$ for any $e=uv\in E(G)$ with $D(e)=(u,v)$ if and only if the function $f$ satisfies the condition $\sum_{e\in E(C)}\sigma_{D(C)}(e)f(e)=0$ for every cycle $C$ of $G$. This condition is equivalent to $\alpha(C)=0$ for all cycles $C$ of $G$ since $\alpha=\alpha_{D,f}$. Therefore, $C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}=0$ whenever $G\langle S\rangle$ is an $\alpha$-incompatible subgraph of $G$. Then, $P\big{(}D(G),f;\Gamma\big{)}$ in (2.2) can be simplified to the following form
$P\big{(}D(G),f;\Gamma\big{)}=\sum_{\begin{subarray}{c}S\subseteq E(G),\\ G\langle S\rangle\mbox{ is $\alpha_{G\langle S\rangle}$-compatible}\end{subarray}}(-1)^{|S|}C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}.$ (2.3)
Denote by $C_{S}\big{(}D\langle S\rangle,f_{S};\Gamma\big{)}$ the collection of vertex colorings $c:V(G)\to\Gamma$ of $G\langle S\rangle$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$. Then, for a fixed vertex coloring $c\in C_{S}\big{(}D\langle S\rangle,f_{S};\Gamma\big{)}$, every vertex coloring $c^{\prime}\in C_{S}\big{(}D\langle S\rangle,f_{S};\Gamma\big{)}$ can gives rise to a unique vertex coloring $c^{\prime}-c\in C_{S}\big{(}D\langle S\rangle,0_{S};\Gamma\big{)}$. Conversely, every vertex coloring $c^{\prime\prime}\in C_{S}\big{(}D\langle S\rangle,0_{S};\Gamma\big{)}$ also gives rise to a unique vertex coloring $c+c^{\prime\prime}\in C_{S}\big{(}D\langle S\rangle,f_{S};\Gamma\big{)}$. Notice
$C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}=|C_{S}\big{(}D\langle S\rangle,f_{S};\Gamma\big{)}|\quad\mbox{\rm~{}and~{}}\quad C_{S}\big{(}D\langle S\rangle,0;\Gamma\big{)}=|C_{S}\big{(}D\langle S\rangle,0_{S};\Gamma\big{)}|.$
It follows that
$C_{S}\big{(}D\langle S\rangle,f;\Gamma\big{)}=C_{S}\big{(}D\langle S\rangle,0;\Gamma\big{)}=|\Gamma|^{c(G\langle S\rangle)}.$
Together with (2.3), we have
$P\big{(}D(G),f;\Gamma\big{)}=\sum_{\begin{subarray}{c}S\subseteq E(G),\\ G\langle S\rangle\mbox{ is $\alpha_{G\langle S\rangle}$-compatible}\end{subarray}}(-1)^{|S|}|\Gamma|^{c(G\langle S\rangle)},$
which completes the proof. ∎
As a direct consequence of Theorem 2.1, the cycle-assigning polynomials corresponding to admissible assignings precisely count the number of $(\Gamma,f)$-colorings of $G$ for any Abelian group $\Gamma$ and function $f:E(G)\to\Gamma$.
Theorem 2.2.
Let $\alpha$ be an admissible assigning of $G$. Then $P(G,\alpha;k)=P\big{(}D(G),f;\Gamma\big{)}$ for every orientation $D\in\mathcal{O}(G)$, every Abelian group $\Gamma$ of order $k$ and every function $f:E(G)\to\Gamma$ satisfying $\alpha_{D,f}=\alpha$.
By applying Theorem 2.2, we obtain the next corollary, which states that the number of $(G,f)$-colorings depends only on $\alpha_{D,f}$, and not on the structure of $\Gamma$ and the choice of the orientation $D$.
[…]
Let $D,D^{\prime}\in\mathcal{O}(G)$, $\Gamma,\Gamma^{\prime}$ be two Abelian groups of the same order and functions $f:E(G)\to\Gamma$, $f^{\prime}:E(G)\to\Gamma^{\prime}$. If $\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$, then $P\big{(}D(G),f;\Gamma\big{)}=P\big{(}D^{\prime}(G),f^{\prime};\Gamma^{\prime}\big{)}$.
Proof.
By Theorem 2.2, we have
$P\big{(}D(G),f;\Gamma\big{)}=P(G,\alpha;k)=P\big{(}D^{\prime}(G),f^{\prime};\Gamma^{\prime}\big{)},$
where $\alpha=\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$. ∎
We further explore the fundamental properties of cycle-assigning polynomials. Suppose $e\in E(G)$ is not a loop of $G$. Then the set $\mathcal{C}(G/e)$ of cycles of $G/e$ consists of the following two parts:
$\mathcal{C}_{-e}(G):=\big{\{}C\in\mathcal{C}(G/e):C\in\mathcal{C}(G),e\notin E(C)\big{\}},\;\mathcal{C}_{/e}(G/e):=\big{\{}C/e:C\in\mathcal{C}(G),e\in E(C)\big{\}}.$
In this sense, each assigning $\alpha$ of $G$ induces an assigning $\alpha_{G/e}$ of $G/e$, where $\alpha_{G/e}(C):=\alpha(C)$ for $C\in\mathcal{C}_{-e}(G)$ and $\alpha_{G/e}(C/e):=\alpha(C)$ for $C\in\mathcal{C}(G)$ containing $e$. We first obtain a deletion-contraction formula for general cycle-assigning polynomials, which provides an inductive approach to proving the Generalized Whitney’s Broken Cycle Theorem (see Theorem 3.1).
Theorem 2.3 (Deletion-contraction Formula).
Let $\alpha$ be an assigning of $G$. Then
$P(G,\alpha;k)=\begin{cases}\frac{k-1}{k}P(G-e,\alpha_{G-e};k),&\mbox{ if }e\mbox{ is a bridge};\\ 0,&\mbox{ if }e\mbox{ is a loop and }\alpha(e)=0;\\ P(G-e,\alpha_{G-e};k),&\mbox{ if }e\mbox{ is a loop and }\alpha(e)=1;\\ P(G-e,\alpha_{G-e};k)-P(G/e,\alpha_{G/e};k),&\mbox{ otherwise}.\end{cases}$
Proof.
Let $\mathcal{S}_{e}$ be the family of $\alpha$-compatible spanning subgraphs of $G$ containing $e$ and $\mathcal{S}_{-e}$ be the family of $\alpha$-compatible spanning subgraphs of $G$ that do not contain $e$. If $e$ is a bridge, then a subgraph $H$ containing $e$ is an $\alpha$-compatible spanning subgraph of $G$ if and only if $H-e$ is an $\alpha$-compatible spanning subgraph of $G$. Therefore, we conclude that every $\alpha$-compatible spanning subgraph $H$ of $G$ containing $e$ gives rise to a unique $\alpha$-compatible spanning subgraph $H-e$ of $G$ that does not contain $e$, and vice versa. Then we have
$P(G,\alpha;k)=\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H+e)|}k^{c(H+e)}+\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H)|}k^{c(H)}.$
Since $e$ is a bridge, $c(H+e)=c(H)-1$ for any $H\in\mathcal{S}_{-e}$. Hence, we deduce
$P(G,\alpha;k)=\frac{k-1}{k}P(G-e,\alpha_{G-e};k).$
If $e$ is a loop, we need to consider both cases: $\alpha(e)=0$ and $\alpha(e)=1$. For the former case, similarly to the case where $e$ is a bridge, we can easily derive $P(G,\alpha;k)=0$. For the latter case, it is clear that $H$ is an $\alpha$-compatible spanning subgraph of $G$ if and only if $H$ is an $\alpha_{G-e}$-compatible spanning subgraph of $G-e$. So, we arrive at $P(G,\alpha;k)=P(G-e,\alpha_{G-e};k)$.
If $e$ is neither a bridge nor a loop, then $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing no $e$ if and only if $H$ is an $\alpha_{G-e}$-compatible spanning subgraph of $G-e$, and $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing $e$ if and only if $H/e$ is an $\alpha_{G/e}$-compatible spanning subgraph of $G/e$. Therefore, we have
$P(G,\alpha;k)$ $=\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H)|}k^{c(H)}+\sum_{H\in\mathcal{S}_{e}}(-1)^{|E(H)|}k^{c(H)}$
$=P(G-e,\alpha_{G-e};k)+\sum_{H\in\mathcal{S}_{e}}(-1)^{|E(H/e)|+1}k^{c(H/e)}$
$=P(G-e,\alpha_{G-e};k)-P(G/e,\alpha_{G/e};k),$
which finishes the proof. ∎
In addition, the cycle-assigning polynomial $P(G,\alpha;k)$ has the following simple decomposition formula.
Theorem 2.4.
Let $\alpha$ be an assigning of $G$ and $G=G_{1}\sqcup G_{2}\sqcup\cdots\sqcup G_{i}$, where $G_{j}$ is the component of $G$ for $j=1,2,\ldots,i$. Then
$P(G,\alpha;k)=\prod_{j=1}^{i}P(G_{j},\alpha_{G_{j}};k).$
Proof.
Since $G$ decomposes into the union of the components $G_{j}$ of $G$, each $\alpha$-compatible spanning subgraph of $G$ is partitioned into the disjoint union of the $\alpha_{G_{j}}$-compatible spanning subgraph of $G_{j}$, and vice versa. Therefore, we can deduce $P(G,\alpha;k)=\prod_{j=1}^{i}P(G_{j},\alpha_{G_{j}};k)$, which finishes the proof. ∎
At the end of the section, we list the cycle-assigning polynomials for the empty graph, a tree and the cycle on $n$ vertices.
[…]
  • (1)
    Let $O_{n}$ be the empty graph on $n$ vertices. Then $\alpha=0$ and
    $P(O_{n},0;k)=k^{n}.$
  • (2)
    Let $T_{n}$ be a tree on $n$ vertices. Then $\alpha=0$ and
    $P(T_{n},0;k)=k(k-1)^{n-1}=\sum_{i=1}^{n}(-1)^{n-i}\binom{n-1}{i-1}k^{i}.$
  • (3)
    Let $C_{n}$ be a cycle on $n$ vertices. Combining part (2) and Theorem 2.3, we have
    $P(C_{n},\alpha;k)=\begin{cases}(k-1)^{n}+(-1)^{n}(k-1),&\mbox{ if }\alpha(C_{n})=0;\\ (k-1)^{n}+(-1)^{n+1},&\mbox{ if }\alpha(C_{n})=1.\end{cases}$
3 Generalization of Whitney’s Broken Cycle Theorem
A natural question is whether there is a combinatorial interpretation for the coefficients of cycle-assigning polynomials. To address this, we introduce the concept of an $\alpha$-compatible broken cycle as an extension of the usual broken cycle.
[…]
Let $\alpha$ be an assigning of $G$ and $\prec$ be a total order on $E(G)$. We say that a subgraph $H$ of $G$ is said to be:
  • an $\alpha$-compatible cycle of $G$ if $H$ is a cycle and $\alpha(H)=0$;
  • an $\alpha$-compatible broken cycle of $G$ if $H$ is obtained from an $\alpha$-compatible cycle by removing the maximal element with respect to the total order $\prec$.
The following result expresses the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ in terms of $\alpha$-compatible spanning subgraphs that do not contain $\alpha$-compatible broken cycles, which is an extended version of the Whitney’s celebrated Broken Cycle Theorem given by Whitney [26]. It is worth noting that the coefficients $w_{i}(G,\alpha)$ in Theorem 3.1 are independent of the choice of edge order.
Theorem 3.1 (Generalized Whitney’s Broken Cycle Theorem).
Let $\alpha$ be an assigning of $G$ and $\prec$ a total order on $E(G)$. Write $P(G,\alpha;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G,\alpha)k^{|V(G)|-i}$, where $r(G)=|V(G)|-c(G)$. Then each unsigned coefficient $w_{i}(G,\alpha)$ is the number of $\alpha$-compatible spanning subgraphs of $G$ that have exactly $i$ edges and do not contain $\alpha$-compatible broken cycles with respect to $\prec$.
Proof.
Let $E(G)=\{e_{1}\prec e_{2}\prec\cdots\prec e_{m}\}$. We prove the result by induction on $m$. If $m=0$, there is nothing to prove. Suppose $m\geq 1$ and the result holds for smaller values of $m$. Let $e_{m}=uv$. If $e_{m}$ is a loop and $\alpha(e_{m})=0$, it is trivial via Theorem 2.3. If $e_{m}$ is a loop and $\alpha(e_{m})=1$, from
$P(G,\alpha;k)=P(G-e,\alpha_{G-e};k)$
in Theorem 2.3, the result holds in this case by the inductive hypothesis.
Finally, if $e_{m}$ is neither a bridge nor a loop, by the inductive hypothesis, the result holds for both graphs $G-e_{m}$ and $G/e_{m}$. With a slight abuse of notation, we identify $E(G/e_{m})=\{e_{1}\prec e_{2}\prec\cdots\prec e_{m-1}\}$. More specifically, by identifying the ends of $e_{m}$, this create a single new vertex of $G/e_{m}$, denoted by $v_{e_{m}}$. Suppose the edge $e_{i}=xy$ in $G$ for some $i<m$. If neither $x$ nor $y$ is one of the vertices $u$ or $v$, then $e_{i}$ automatically remains an edge in $G/e_{m}$. If either $x$ or $y$ is one of the vertices $u$ or $v$, we identity $e_{i}$ as $uv_{e_{m}}$, $v_{e_{m}}v$ or $v_{e_{m}}v_{e_{m}}$ depending on the incidence of $e_{i}$ in $G$.
By the Deletion-contraction Formula in Theorem 2.3, we have $w_{0}(G,\alpha)=w_{0}(G-e_{m},\alpha_{G-e_{m}})$ and
$w_{i}(G,\alpha)=w_{i}(G-e_{m},\alpha_{G-e_{m}})+w_{i-1}(G/e_{m},\alpha_{G/e_{m}}),\;i=1,2,\ldots,r(G).$ (3.1)
Therefore, to complete the induction step, it is sufficient to check that the relation in (3.1) is true. But this is a consequence of the following two simple facts: (a) $H$ is an $\alpha$-compatible spanning subgraph $H$ of $G$ that does not contain $e$ and $\alpha$-compatible broken cycles if and only if $H$ is an $\alpha_{G-e_{m}}$-compatible spanning subgraph of $G-e_{m}$ that does not contain $\alpha_{G-e_{m}}$-compatible broken cycles; (b) $H$ is an $\alpha$-compatible spanning subgraph $H$ of $G$ that contains $e$ and no $\alpha$-compatible broken cycles if and only if $H/e_{m}$ is an $\alpha_{G/e_{m}}$-compatible spanning subgraph of $G/e_{m}$ that does not contain $\alpha_{G/e_{m}}$-compatible broken cycles. So, we complete the proof. ∎
In addition, when $\alpha$ is an admissible assigning of $G$, we provide another direct proof of Theorem 3.1. For this purpose, the next lemma is required.
[…]
Let $\alpha$ be a $D$-admissible assigning of $G$ and $\prec$ a total order on $E(G)$. Suppose $H$ is a subgraph of $G$ containing an $\alpha$-compatible broken cycle $B$ and $B=C-e_{C}$ for some $\alpha$-compatible cycle $C$ of $G$, where $e_{C}$ is the maximal member in $E(C)$ with respect to $\prec$. If $H$ is an $\alpha$-compatible subgraph of $G$ and $e_{C}\notin E(H)$, then $H+e_{C}$ is also an $\alpha$-compatible subgraph of $G$.
Proof.
Note that $C$ is a cycle in $H+e_{C}$, but not in $H$. If $H+e_{C}$ contains exactly one more cycle $C$ than $H$, there is nothing left to prove. Otherwise, $H+e_{C}$ contains at least one additional cycle compared to $H$, excluding the cycle $C$. Notably, every such additional cycle, together with the cycle $C$, forms a theta-subgraph in $H+e_{C}$. Therefore, this case reduces to verifying that each such the theta-subgraph in $H+e_{C}$ is an $\alpha$-compatible subgraph of $G$. Notice that $\alpha$ is a $D$-admissible assigning of $G$. Without loss of generality, we only need to examine the following oriented theta-subgraph in $H+e_{C}$ with the orientation $D$, as shown in Figure 1.
[…]
Figure 1: The directed paths $P_{1},P_{2}$ and $P_{3}$ toghter form an oriented theta-graph
More specifically, in Figure 1, the edge $e_{C}=u_{0}v_{0}$; the cycle $C$, which is the union of the red directed path $P_{1}=(u_{i},u_{i-1},\ldots,v_{k-1},v_{k})$ and the green directed path $P_{2}=(v_{k},v_{k+1},\ldots,v_{l},u_{i})$, and the cycle $C_{1}$, which is the union of the red directed path $P_{1}$ and the blue directed path $P_{3}=(u_{i},u_{i+1},\ldots,u_{j},v_{k})$, are in $H+e_{C}$ but not in $H$; the cycle $C_{2}$ is the common cycle of $H+e_{C}$ and $H$, formed by the union of both directed paths $P_{2}$ and $P_{3}$. Then, to complete our proof, it suffices to show $\alpha(C)=\alpha(C_{1})=\alpha(C_{2})=0$.
Since $\alpha$ is $D$-admissible, we may assume $\alpha=\alpha_{D,f}$ for some Abelian group $\Gamma$ and function $f:E(G)\to\Gamma$. Combining that $C$ and $H$ are $\alpha$-compatible subgraphs of $G$, this means
[…]
This is further equivalent to satisfying the following equations
$\sum_{e\in E(C)}\sigma_{D(C)}(e)f(e)=\sum_{e\in E(P_{1})}\sigma_{D(C)}(e)f(e)+\sum_{e\in E(P_{2})}\sigma_{D(C)}(e)f(e)=0$
and
$\sum_{e\in E(C_{2})}\sigma_{D(C_{2})}(e)f(e)=\sum_{e\in E(P_{2})}\sigma_{D(C_{2})}(e)f(e)+\sum_{e\in E(P_{3})}\sigma_{D(C_{2})}(e)f(e)=0.$
Together with $\sigma_{D(C_{1})}(e)=-\sigma_{D(C)}(e)$ for any edge $e\in E(P_{1})$ and $\sigma_{D(C_{1})}(e)=\sigma_{D(C_{2})}(e)$ for any edge $e\in E(P_{3})$, we deduce
$\sum_{e\in E(C_{1})}\sigma_{D(C_{1})}(e)f(e)=\sum_{e\in E(P_{1})}\sigma_{D(C_{1})}(e)f(e)+\sum_{e\in E(P_{3})}\sigma_{D(C_{1})}(e)f(e)=0.$
This implies $\alpha(C_{1})=\alpha_{D,f}(C_{1})=0$. Therefore, we conclude that $H+e_{C}$ is an $\alpha$-compatible subgraph of $G$, which finishes the proof. ∎
We now return to the proof of the Generalized Whitney’s Broken Cycle Theorem when $\alpha$ is an admissible assigning.
Proof of Theorem 3.1.
Suppose $G$ has $q\geq 0$ $\alpha$-compatible broken cycles. For every $\alpha$-compatible broken cycle $B$, let $e_{B}$ be the maximal member in $B$ with respect to the total order $\prec$. We can arrange the $\alpha$-compatible broken cycles $B_{1},\ldots,B_{q}$ such that $e_{B_{1}}\preceq e_{B_{2}}\cdots\preceq e_{B_{q}}$. Let $\mathcal{S}_{1}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing the $\alpha$-compatible broken cycle $B_{1}$. For $i=2,\ldots,q$, let $\mathcal{S}_{i}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing the $\alpha$-compatible broken cycle $B_{i}$ but not any of $B_{1},\ldots,B_{i-1}$. In addition, let $\mathcal{S}_{q+1}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing no $\alpha$-compatible broken cycles. Then $\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{q+1}\}$ decomposes the set of $\alpha$-compatible spanning subgraphs of $G$. Combining Theorem 2.2, the cycle-assigning polynomial $P(G,\alpha;k)$ can be written in the form
$P(G,\alpha;k)=\sum_{i=1}^{q+1}\sum_{H\in\mathcal{S}_{i}}(-1)^{|E(H)|}k^{c(H)}.$ (3.2)
We first consider $\mathcal{S}_{1}$. Assume $B_{1}=C_{1}-e_{C_{1}}$ for some $\alpha$-compatible cycle $C_{1}$ of $G$, where $e_{C_{1}}$ is the maximal member in $C_{1}$ with respect to $\prec$. Clearly each member $H$ in $\mathcal{S}_{1}$ containing $e_{C_{1}}$ gives rise to a unique member $H-e_{C_{1}}$ in $\mathcal{S}_{1}$ that does not contain $e_{C_{1}}$. Conversely, according to Section 3, each member $H$ in $\mathcal{S}_{1}$ containing no $e_{C_{1}}$ gives rise to a unique member $H+e_{C_{1}}$ in $\mathcal{S}_{1}$ containing $e_{C_{1}}$. In addition, since $B_{1}+e_{C_{1}}=C_{1}$ is a cycle of $G$, for each $H\in\mathcal{S}_{1}$ containing $e_{C_{1}}$, we have $c(H)=c(H-e_{C_{1}})$. Thus, we derive
$\sum_{H\in\mathcal{S}_{1},\,e_{C_{1}}\in E(H)}\big{(}(-1)^{|E(H)|}k^{c(H)}+(-1)^{|E(H)|-1}k^{c(H-e_{C_{1}})}\big{)}=0.$
Namely, we arrive at $\sum_{H\in\mathcal{S}_{1}}(-1)^{|E(H)|}k^{c(H)}=0$ in (3.2).
Consider $\mathcal{S}_{2}$. Assume $B_{2}=C_{2}-e_{C_{2}}$ for some $\alpha$-compatible cycle $C_{2}$ of $G$, where $e_{C_{2}}$ is the maximal member in $C_{2}$ with respect to $\prec$. It follows from the ordering $e_{B_{1}}\preceq e_{B_{2}}$ that $e_{B_{1}}\preceq e_{B_{2}}\prec e_{C_{2}}$. So, $B_{1}$ and $B_{2}$ do not contain $e_{C_{2}}$. Combining this, using the same argument as the case $\mathcal{S}_{1}$, we conclude that every element $H$ in $\mathcal{S}_{2}$ containing no $e_{C_{2}}$ gives rise to a unique member $H+e_{C_{2}}$ in $\mathcal{S}_{2}$ containing $e_{C_{2}}$, and vice versa. Hence, we also obtain $\sum_{H\in\mathcal{S}_{2}}(-1)^{|E(H)|}k^{c(H)}=0$ in (3.2).
Likewise, we can deduce the total contribution of the members in $\mathcal{S}_{i}$ to $P(G,\alpha;k)$ is zero for each $i=1,2,\ldots,q$. Therefore, $P(G,\alpha;k)$ in (3.2) can be written as
$P(G,\alpha;k)=\sum_{H\in\mathcal{S}_{q+1}}(-1)^{|E(H)|}k^{c(H)}.$ (3.3)
Since every element $H\in\mathcal{S}_{q+1}$ is an $\alpha$-compatible spanning subgraph of $G$ and contains no $\alpha$-compatible broken cycles, we can deduce that $H$ contains no cycles. Therefore, $H$ is a forest. This implies that for any member $H\in\mathcal{S}_{q+1}$,
$|E(H)|\leq r(G)\quad\mbox{\rm~{}and~{}}\quad c(H)=|V(G)|-|E(H)|.$
Thus, $P(G,\alpha;k)$ in (3.3) can be further simplified to the following form
$P(G,\alpha;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G,\alpha)k^{|V(G)|-i},$
which finishes the proof. ∎
Furthermore, as a direct result of Theorem 3.1, we have the next corollary.
[…]
Let $\alpha$ be an assigning of $G$. Then $(-1)^{|V(G)|}P(G,\alpha;-1)$ is the number of $\alpha$-compatible spanning subgraphs of $G$ that do not contain $\alpha$-compatible broken cycles with respect to a total order $\prec$.
Notably, every $(\Gamma,0)$-coloring of $G$ is exactly a proper coloring of $G$, vice versa. Therefore, $P(G,0;k)$ is precisely the classical chromatic polynomal $P\big{(}G;k\big{)}$ of $G$. In this case, the 00-compatible spanning subgraph, 00-compatible cycle and 00-compatible broken cycle of $G$ correspond to the ordinary spanning subgraph, cycle and broken cycle of $G$, respectively. Hence, Theorem 3.1 deduces the following Whitney’s Broken Cycle Theorem in [26] directly.
Theorem 3.2 ([26]).
Let $\prec$ be a total order on $E(G)$. Write the chromatic polynomial $P(G;k)$ as $P(G;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G)k^{|V(G)|-i}$. Then each unsigned coefficient $w_{i}(G)$ is the number of spanning subgraphs of $G$ that have exactly $i$ edges and do not contain broken cycles with respect to the total order $\prec$.
4 Coefficient comparisons and further explorations
In this section, we primarily apply the Generalized Broken Cycle Theorem to explore how the coefficients of cycle-assigning polynomials vary with changes in assignings and to handle certain group coloring problems in graphs. We first establish a unified order-preserving relation from assignings to cycle-assigning polynomials when both are naturally ordered. Let $\alpha$ be an assigning of $G$ and $\prec$ be a total order on $E(G)$. For convenience, we denote by $\mathcal{C}(G,\alpha)$ and $\mathcal{C}^{\prime}(G,\alpha)$ the sets of $\alpha$-compatible cycles and $\alpha$-compatible broken cycles of $G$ with respect to $\prec$, respectively. Additionally, let $\mathcal{C}^{\prime\prime}(G,\alpha)$ represent the set of $\alpha$-compatible spanning subgraphs of $G$ that do not contain $\alpha$-compatible broken cycles with respect to $\prec$.
Theorem 4.1 (Comparison of Coefficients).
Let $\alpha,\alpha^{\prime}$ be two assignings of $G$. If $\alpha(C)\leq\alpha^{\prime}(C)$ for all cycles $C\in\mathcal{C}(G)$, then the unsigned coefficients of the cycle-assigning polynomials $P(G,\alpha;k)$ and $P(G,\alpha^{\prime};k)$ satisfy
$w_{i}(G,\alpha)\leq w_{i}(G,\alpha^{\prime})\quad\mbox{ for }\quad i=0,1,\ldots,r(G).$
In particular, $w_{i}(G)=w_{i}(G,0)\leq w_{i}(G,\alpha)$ for any assigning $\alpha$ of $G$.
Proof.
By Theorem 3.1, the proof reduces to showing that every member $H$ of $\mathcal{C}^{\prime\prime}(G,\alpha)$ belongs to $\mathcal{C}^{\prime\prime}(G,\alpha^{\prime})$. Given an element $H\in\mathcal{C}^{\prime\prime}(G,\alpha)$. As $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing no $\alpha$-compatible broken cycles, this implies that $H$ does not contain cycles. Furthermore, since $\alpha(C)\leq\alpha^{\prime}(C)$ for any cycle of $G$, we have $\mathcal{C}(G,\alpha)\supseteq\mathcal{C}(G,\alpha^{\prime})$. This means $\mathcal{C}^{\prime}(G,\alpha)\supseteq\mathcal{C}^{\prime}(G,\alpha^{\prime})$. This further deduces that $H$ does not contain $\alpha^{\prime}$-compatible broken cycles. Therefore, we conclude that $H$ is an $\alpha^{\prime}$-compatible spanning subgraph of $G$ containing no $\alpha^{\prime}$-compatible broken cycles. Namely, $\mathcal{C}^{\prime\prime}(G,\alpha)\subseteq\mathcal{C}^{\prime\prime}(G,\alpha^{\prime})$. ∎
When $G$ is a loopless graph, each unsigned coefficient $w_{i}(G)$ of $P(G;k)$ is a positive integer, which is easily obtained from Theorem 3.2. Applying Theorem 4.1 to this, we can derive that for any loopless graph $G$, the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ are nonzero and alternate in sign.
[…]
Let $\alpha$ be an assigining of $G$. If $G$ is a loopless graph, then each unsigned coefficient $w_{i}(G,\alpha)$ of the cycle-assigning polynomial $P(G,\alpha;k)$ is a positive integer. Moreover, $w_{0}(G,\alpha)=1$.
It is well-known that the unimodality of the sequence $w_{i}(G)$ was conjectured by Read [23], and the log-concavity was conjectured by Hoggar [9]. Both of these conjectures have been confirmed by Huh et al. [1, 10, 11]. Inspired by their work, we conjecture that the unsigned coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ form a log-concave, and hence unimodal, sequence for all loopless graphs $G$.
[…]
Let $G$ be a loopless graph and $\alpha$ be an assigning of $G$. Then the sequence $w_{0}(G,\alpha),w_{1}(G,\alpha),\ldots,w_{r(G)}(G,\alpha)$ satisfies the following properties:
  • The sequence $w_{i}(G,\alpha)$ is unimodal, i.e., there is an index $1\leq j\leq r(G)$ such that
    $w_{0}(G,\alpha)\leq\cdots\leq w_{j-1}(G,\alpha)\leq w_{j}(G,\alpha)\geq w_{j+1}(G,\alpha)\geq\cdots\geq w_{r(G)}(G,\alpha).$
  • The sequence $w_{i}(G,\alpha)$ is log-concave, i.e, for $j=1,\ldots,r(G)-1$,
    $w_{j-1}(G,\alpha)w_{j+1}(G,\alpha)\leq w_{j}(G,\alpha)^{2}.$
It is a pleasant surprise that the Generalized Broken Cycle Theorem provides a new perspective on addressing certain group coloring problems in graphs. A well-known property in [25] says that if a graph has a nowhere-zero $\Gamma$-flow, then it admits a nowhere-zero $\Gamma^{\prime}$-flow for any Abelian group $\Gamma^{\prime}$ whose order is at least the order of $\Gamma$. This does not extend to the more general concept of group connectivity observed by Jaeger et al. [12]. For a connected planar graph $G$, Jaeger et al. [12] also showed that $G$ is $\Gamma$-colorable if and only if the dual graph $G^{*}$ is $\Gamma$-connected. Naturally, there is a corresponding problem in group colorings: for any Abelian groups $\Gamma_{1}$ and $\Gamma_{2}$ with the same order, is a $\Gamma_{1}$-colorable graph also $\Gamma_{2}$-colorable?
Most recently, Langhede and Thomassen in [19] confirmed that the property holds if $\Gamma^{\prime}$ is sufficiently large compared to $\Gamma$. This further implied that $G$ is always $\Gamma$-colorable whenever $\Gamma$ has a sufficiently large order, which can be explained by the Generalized Broken Cycle Theorem. More specifically, note the basic fact that for a loopless graph $G$, $G$ is $\Gamma$-colorable for some Abelian group $\Gamma$ if and only if $P\big{(}D(G),f;\Gamma\big{)}>0$ for any $f:E(G)\to\Gamma$. This is also equivalent to saying that $G$ is $\Gamma$-colorable if and only if $P(G,\alpha_{G,f};|\Gamma|)>0$ for any $D$-admissible assigning $\alpha_{D,f}$ related to $f:E(G)\to\Gamma$. When $G$ is a loopless graph, it follows from Theorem 3.1 that $w_{0}(G,\alpha)=1$ for any assigning $\alpha$ of $G$. Thus, $P(G,\alpha;|\Gamma|)>0$ holds for all assigning $\alpha$ of $G$ in this case whenever the Abelian group $\Gamma$ has a large enough order.
We conclude this section with the following corollary, which is a direct consequence of the mentioned fact.
[…]
Let $\Gamma,\Gamma^{\prime}$ be two Abelian groups of the same order, $G$ be a $\Gamma^{\prime}$-colorable graph and $D,D^{\prime}\in\mathcal{O}(G)$. If each $D$-admissible assigning $\alpha_{D,f}$ of $G$ with $f:E(G)\to\Gamma$ is also a $D^{\prime}$-admissible assigning $\alpha_{D^{\prime},f^{\prime}}$ of $G$ for some $f^{\prime}:E(G)\to\Gamma^{\prime}$, then $G$ is $\Gamma$-colorable.
Proof.
Since $G$ is $\Gamma^{\prime}$-colorable, we have $P\big{(}D^{\prime}(G),f^{\prime};\Gamma^{\prime}\big{)}>0$ for any $f^{\prime}:E(G)\to\Gamma^{\prime}$. As each $D$-admissible assigning $\alpha_{D,f}$ with $f:E(G)\to\Gamma$ is a $D^{\prime}$-admissible assigning $\alpha_{D^{\prime},f^{\prime}}$ of $G$ for some $f^{\prime}:E(G)\to\Gamma^{\prime}$, it follows that for any $f:E(G)\to\Gamma$,
$P\big{(}D(G),f;\Gamma\big{)}=P(G,\alpha;|\Gamma|)=P\big{(}D^{\prime}(G),f^{\prime};\Gamma^{\prime}\big{)}>0$
via Section 2, where $\alpha=\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$. This implies that $G$ is $\Gamma$-colorable. ∎
Acknowledgements
The author is supported by National Natural Science Foundation of China under Grant No. 12301424.