### v4

# Polynomials Counting Group Colorings in Graphs

Abstract
Jaeger et al. in 1992 introduced group coloring as the dual concept to group connectivity in graphs. Let $A$ be an additive Abelian group, $f:E(G)\to A$ and $D$ an orientation of a graph $G$. A vertex coloring $c:V(G)\to A$ is an $(A,f)$-coloring if $c(v)-c(u)\neq f(e)$ for each oriented edge $e=uv$ from $u$ to $v$ with respect to $D$. Kochol recently introduced the assigning polynomial counting nowhere-zero chains in graphs–nonhomogeneous analogues of nowhere-zero flows in [13], and later extended this approach to regular matroids in [14]. Motivated by Kochol’s work, we define the $\alpha$-compatible graph and the cycle-assigning polynomial $P(G,\alpha;k)$ at $k$ in terms of $\alpha$-compatible spanning subgraphs, where $\alpha$ is an assigning of $G$ from its cycles to $\{0,1\}$. We prove that $P(G,\alpha;k)$ evaluates the number of $(A,f)$-colorings of $G$ for any Abelian group $A$ of order $k$ and $f:E(G)\to A$ such that the assigning $\alpha_{D,f}$ given by $f$ equals $\alpha$. In particular, $P(G,\alpha;k)$ is the chromatic polynomial of $G$ if $\alpha$ is the zero function. Based on Kochol’s work, we also show that $k^{-c(G)}P(G,\alpha;k)$ is a polynomial enumerating $(A,f)$-tensions and counting specific nowhere-zero chains.
By extending the concept of broken cycle to broken compatible cycles, associated with any admissible assigning $\alpha$ of $G$, we demonstrate that the absolute value of the coefficient of $k^{|V(G)|-i}$ in $P(G,\alpha;k)$ equals the number of $\alpha$-compatible spanning subgraphs that have $i$ edges and contain no broken $\alpha$-compatible cycles. This generalizes the Whitney’s celebrated Broken Cycle Theorem. Based on the combinatorial explanation, we establish a unified order-preserving relation from admissible assignings to cycle-assigning polynomials. Finally, we show that for any admissible assigning $\alpha$ of $G$ with $\alpha(e)=1$ for every loop $e$ of $G$, the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ are nonzero and alternate in sign.
Keywords: graph coloring, tension, chromatic polynomial, assigning polynomial, $\alpha$-compatible graph, broken cycle
Mathematics Subject Classifications: 05C31, 05C15
1 Introduction
The graphs $G$ with vertex set $V(G)$ and edge set $E(G)$ considered in this paper are finite, allowing multiple edges and loops. A more general setting and background on graphs can be found in the book [4]. An orientation of a graph is an assignment of direction to each edge. In this context, each oriented edge is referred to as an arc. An oriented graph arises from a graph $G$ together with a particular orientation $D$, denoted by $D(G):=\big{(}V(D),A(D)\big{)}$, where $V(D)=V(G)$ and $A(D)$ is the collection of arcs with respect to the orientation $D$.
Graph coloring, originating from the 19th century Four Color Problem, is a significant subfield in graph theory. A $k$-vertex coloring, or simply a $k$-coloring, of $G$ is a mapping $c:V(G)\to S$, where $S$ is a set of distinct $k$ colors, typically $S=\{1,2,...,k\}$. A vertex coloring is said to be proper if no two adjacent vertices share the same color.
We consider finite additive Abelian groups $A$ with the identity 00 as coloring sets in this paper. Jaeger, Linial, Payan and Tarsi [10] introduced the concept of group coloring as the dual concept of group connectivity of graphs in 1992. Group connectivity and group colorings of graphs are nicely surveyed in [17]. Let $f:E(G)\to A$ be a function. A vertex coloring $c:V(G)\to A$ is called an $(A,f)$-coloring if $c(v)-c(u)\neq f(e)$ for each oriented edge $e=uv$ from $u$ to $v$ with respect to an orientation $D$ of $G$. When $f\equiv 0$, $(A,0)$-coloring is exactly the ordinary proper coloring. Furthermore, $G$ is said to be $A$-colorable under the orientation $D$ if for any function $f:E(G)\to A$, $G$ always admits an $(A,f)$-coloring. Since then, topics related to group coloring problems have attracted a lot of research attentions, such as [15, 21, 22, 23, 16, 18, 20]. Their primary focus is on the (weak) group chromatic number and related group coloring problems, such as whether a graph that is $A_{1}$-colorable for some Abelian group $A_{1}$ is also $A_{2}$-colorable for any Abelian group $A_{2}$ of the same order. However, another fundamental problem of counting $(A,f)$-colorings has been overlooked. In this paper, we introduce the cycle-assigning polynomial, a concept inspired by Kochol’s work [13, 14], which precisely counts the number of $(A,f)$-colorings of graphs and generalizes the classical chromatic polynomial. We primarily investigate its fundamental properties, particularly focusing on its coefficients.
A fundamental problem in graph theory is determining the chromatic number of a graph, which refers to the smallest integer $k$ such that the graph admits a proper coloring with $k$ colors. Counting proper colorings can be seen as another facet of this problem. In an attempt to address the long-standing Four Color Problem, Birkhoff [2] discovered that the number of proper colorings of a planar $G$ with $k$ colors is a polynomial in $k$, known as the chromatic polynomial. Whitney [30, 31] later extended the concept of the chromatic polynomial to general graphs and gave a combinatorial description of its coefficients by introducing the broken cycle, now famously known as the Whitney’s Broken Cycle Theorem. Since then, various generalizations of this theorem have been developed across different mathematical structures, such as matroids [5, 9], lattices [3, 26], hyperplane arrangements [24], hypergraphs [7]. See [8] for more related work. Motivated by these work, it is natural to ask the two questions:
  • Is there a polynomial function of $|A|$ that generalizes the chromatic polynomial and counts $(A,f)$-colorings in graphs?
  • Can we provide a combinatorial interpretation for the coefficients of the polynomial function?
It is worth remarking that the difference between the Abelian group colorings of vertices in an oriented graph gives rise to a group-valued function on arcs, refereed to as a tension, whose value at each arc is the head-point value minus the tail-point value of the coloring. From this perspective, the proper colorings correspond to nowhere-zero tensions. It is known that the number of nowhere-zero tensions in Abelian groups of orders $k$ is also a polynomial. This was first introduced by Tutte [28] and later formally named the tension polynomial of $G$ by Kochol [11]. Note that the tension polynomial is nontrivial divisor of the chromatic polynomial. For more related work, refer to [6, 12]. Likewise, an arbitrary $(A,f)$-coloring of $G$ also produces an $(A,f)$-tension, whose detailed definition is presented in Section 3. Naturally, there are two similar questions related to $(A,f)$-tensions, analogous to those mentioned in $(A,f)$-colorings. The answer to the first question in $(A,f)$-tensions seems to be implicit in the work of Kochol [14]. We shall present a precise answer in Section 3, and further explore the connection between $(A,f)$-colorings and nowhere-zero chains related to homomorphisms (see [14]) in Section 3.
To answer these questions, several significant works should be noted. Most recently, Kochol [13] first introduced the assigning polynomial counting nowhere-zero chains in graphs–nonhomogeneous analogues of nowhere-zero flows (called $(A,b)$-flows in [19]), which is a generalization of the classical flow polynomial [28]. Subsequently, Kochol [14] further extended the approach for a regular matroid to count nowhere-zero chains associated with homomorphisms and zero-one assignings from its circuits to $\{0,1\}$. Kochol’s pioneering work has been a cornerstone of our research. Particularly, the zero-one assignings on cycles of graphs play a central role in the study of $(A,f)$-colorings and $(A,f)$-tensions in graphs.
A cycle $C$ of $G$ is a connected 2-regular subgraph of $G$. The corresponding edge set $E(C)$ is referred to as a circuit of $G$. Let $\mathcal{C}(G)$ be the family of cycles of $G$. An assigning of $G$ is a mapping from $\mathcal{C}(G)$ to the set $\{0,1\}$. By convention, we define $\alpha\equiv 0$ when $G$ does not contain cycles, i.e., $\mathcal{C}(G)=\emptyset$.
To more conveniently investigate the properties of cycle-assigning polynomials, we use the concept of the $\alpha$-compatible graph to mirror the role of $(M,\prec)$-compatible and $\delta(M,\alpha;X)$ associated with a matroid $M$ introduced by [14] (see Section 3 for detailed definitions).
[…]
Let $\alpha$ be an assigning of $G$. A graph $G$ is said to be $\alpha$-compatible if $\alpha(C)=0$ for every cycle $C$ in $G$. Furthermore, a subgraph $H$ of $G$ is called an $\alpha$-compatible subgraph if the restriction of $\alpha$ to $H$ satisfies $\alpha(C)=0$ for all cycles $C$ in $H$. In this sense, $H$ is also $\alpha_{H}$-compatible, where $\alpha_{H}$ is the restriction of $\alpha$ to $H$ defined by $\alpha_{H}(C):=\alpha(C)$ for any cycle $C$ of $H$.
Let $f:E(G)\to A$, and $P\big{(}D(G),f;A\big{)}$ denote the number of $(A,f)$-colorings of $G$ with respect to an orientation $D$. We first provide an explicit formula for the counting function $P\big{(}D(G),f;A\big{)}$ in terms of the $\alpha$-compatible spanning subgraph expansion in Theorem 2.1, where $\alpha$ equals the admissible assigning $\alpha_{D,f}$ induced by $f$. The detailed definition of an admissible assigning can be found at the end of the second paragraph in Section 2. Inspired by this expression, we define a cycle-assigning polynomial in terms of $\alpha$-compatible spanning subgraphs. It turns out that the cycle-assigning polynomials precisely count $(A,f)$-colorings of graphs, as shown in Theorem 2.2.
[…]
For any assigning $\alpha$ of $G$, the cycle-assigning polynomial of $G$ in $k$ is defined as
$P(G,\alpha;k):=\sum_{\begin{subarray}{c}\mbox{$H$ is an $\alpha$-compatible}\\ \mbox{spanning subgraph of $G$}\end{subarray}}(-1)^{|E(H)|}k^{c(H)},$
where $c(H)$ is the number of components of $H$.
Furthermore, let
$\tau(G,\alpha;k):=\sum_{\begin{subarray}{c}\mbox{$H$ is an $\alpha$-compatible}\\ \mbox{spanning subgraph of $G$}\end{subarray}}(-1)^{|E(H)|}k^{r(G)-r(H)},$
where $r(H):=|V(G)|-c(H)$. To better distinguish both polynomials $P(G,\alpha;k)$ and $\tau(G,\alpha;k)$, we call $\tau(G,\alpha;k)$ an $\alpha$-assigning polynomial of $G$ originating from [13, 14]. It should be noted that the $(A,f)$-tension serves as a bridge connecting $(A,f)$-colorings and nowhere-zero chains associated with regular matroids and homomorphisms (see Section 3 for detailed discussion).
Due to Kochol’s work [13, 14], we show that $\tau(G,\alpha;k)$ not only enumerates $(A,f)$-tensions of $G$ but also counts the number of certain nowhere-zero chains in $A^{E(G)}$ associated with a specific homomorphism from the regular chain group $\Gamma(G)$ of the cycle matroid $M(G)$ to $A$, as stated in Section 3. As a direct consequence of [14, Theorem 2], an alternative expression for $P(G,\alpha;k)$ and $\tau(G,\alpha;k)$ is presented in Section 3.
To address the second question, we further introduce the concept of a broken $\alpha$-compatible cycle as an extension of the usual broken cycle proposed by Whitney [30]. In Theorem 4.1, when $\alpha$ is an admissible assigning of $G$, we demonstrate that the unsigned coefficients of cycle-assigning polynomials can be expressed in terms of $\alpha$-compatible spanning subgraphs that do not contain broken $\alpha$-compatible cycles. This extends the Whitney’s celebrated Broken Cycle Theorem [30]. Based on the counting interpretation, we further establish a unified order-preserving relation from admissible assignings to cycle-assigning polynomials when both are naturally ordered in Theorem 5.1. In addition, we also provide a new perspective on addressing certain group coloring problems in graphs. A detailed discussion will appear at the end of Section 5.
Notice that the $\alpha$-assigning polynomial $\tau(G,\alpha;k)$ is the nontrivial divisor of the cycle-assigning polynomial $P(G,\alpha;k)$, that is,
$P(G,\alpha;k)=k^{c(G)}\tau(G,\alpha;k).$
Therefore, the coefficients of $\tau(G,\alpha;k)$ share the same properties as those of $P(G,\alpha;k)$.
The paper is organized as follows. Section 2 presents an explicit formula for calculating $(A,f)$-colorings (see Theorem 2.1), and further shows that the number of $(A,f)$-colorings with the function $f$ corresponding to an assigning is evaluated by the cycle-assigning polynomial of $|A|$ in Theorem 2.2. In Section 3, we mainly explain the close relation between the cycle-assigning polynomial and Kochol’s work [14]. In Section 4, associated with an admissible assigning $\alpha$, we interpret $P(G,\alpha;k)$ in terms of $\alpha$-compatible spanning subgraphs that do not contain broken $\alpha$-compatible cycles in Theorem 4.1. Section 5 utilizes the counting interpretation to establish a unified order-preserving relation from admissible assignings to cycle-assigning polynomials in Theorem 5.1, and also demonstrates that for any graph $G$ and its admissible assigning $\alpha$ with $\alpha(e)=1$ for every loop $e$ of $G$, the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$ are nonzero and alternate in sign Section 5.
2 Cycle-assigning polynomials
In this section, we focus on the study of cycle-assigning polynomials, exploring their structural properties and their role in counting $(A,f)$-colorings. Let us review some necessary notations and definitions on graphs. Throughout this paper, we always denote by $S^{E}$ the collection of mappings from a finite set $E$ to a set $S$, unless otherwise stated. The elements of $S^{E}$ are considered as vectors indexed by $E$. An edge with identical ends is called a loop, and an edge with distinct ends is a link. A graph $H$ is a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. When $V(H)=V(G)$, the subgraph $H$ of $G$ is referred to as a spanning subgraph. For an edge subset $S\subseteq E(G)$, $G-S$ is the subgraph obtained by removing the edges in $S$. Denote by $G|S$ the spanning subgraph with edges $S$, that is, $G|S=G-(E(G)\smallsetminus S)$. Conversely, suppose $H$ is a proper subgraph containing no edge $e\in E(G)$, we denote by $H+e$ the subgraph of $G$ by adding the edge $e$. In addition, if $e$ is a link of $G$, the graph $G/e$ is obtained by contracting $e$, that is, identifying the ends of $e$ to create a single new vertex and then removing $e$.
Unless otherwise stated, we always choose an arbitrary but fixed reference orientation $D$ for $G$. For every cycle $C$ of $G$, the orientation $D$ of $G$ gives rise to a map $\eta_{D(C)}$ from the edge set $E(G)$ to the set $\{0,\pm 1\}$ defined by
$\eta_{D(C)}(e):=\begin{cases}1,&\mbox{if $e\in E(C)$, the direction of $D(e)$ is clockwise with respect to $C$};\\ -1,&\mbox{if $e\in E(C)$, the direction of $D(e)$ is counterclockwise with respect to $C$};\\ 0,&\mbox{if $e\in E(G)\smallsetminus E(C)$},\end{cases}$ (2.1)
where $D(e)$ denotes the corresponding arc of the edge $e$ with respect to $D$. Remarkably, each function $f:E(G)\to A$ automatically induces an assigning $\alpha_{D,f}$ on $\mathcal{C}(G)$ such that for each cycle $C$ of $G$,
$\alpha_{D,f}(C)=\begin{cases}0,&\mbox{ if }\sum_{e\in E(C)}\eta_{D(C)}(e)f(e)=0;\\ 1,&\mbox{ otherwise }.\end{cases}$
We say that $\alpha=\alpha_{D,f}$ is $D$-admissible, or simply admissible.
Recall from Section 1 that the counting function $P\big{(}D(G),f;A\big{)}$ is the number of $(A,f)$-colorings of $G$ associated with the orientation $D$ and a function $f:E(G)\to A$. To derive an explicit formula for $P\big{(}D(G),f;A\big{)}$, the Möbius Inversion Formula in [4, Theorem 2.25] is required. It states that if the function $g:2^{E(G)}\to A$ is defined as
$g(S):=\sum_{S\subseteq X\subseteq E(G)}h(X)\quad\mbox{ for }\quad\forall\,S\subseteq E(G),$
then, for all $S\subseteq E(G)$,
$h(S)=\sum_{S\subseteq X\subseteq E(G)}(-1)^{|X|-|S|}g(X),$ (2.2)
where $h$ is a function from the family $2^{E(G)}$ of all edge subsets of $G$ to $A$.
The following result shows that the counting function $P\big{(}D(G),f;A\big{)}$ can be expressed in terms of the $\alpha_{D,f}$-compatible spanning subgraph expansion.
Theorem 2.1 (Counting Formula).
Let $f:E(G)\to A$ and $\alpha=\alpha_{D,f}$. Then
$P\big{(}D(G),f;A\big{)}=\sum_{\begin{subarray}{c}\mbox{$H$ is an $\alpha$-compatible}\\ \mbox{spanning subgraph of $G$}\end{subarray}}(-1)^{|E(H)|}|A|^{c(H)}.$
Proof.
For any $S\subseteq E(G)$, let $C_{S}\big{(}D(G),f;A\big{)}$ be the number of vertex colorings $c:V(G)\to A$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$, and $P_{S}\big{(}D(G),f;A\big{)}$ be the number of vertex colorings $c:V(G)\to A$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$ and $c(v)-c(u)\neq f(e)$ for each edge $e=uv\in E(G)\smallsetminus S$ with $D(e)=(u,v)$. Then we have
$C_{S}\big{(}D(G)|S,f;A\big{)}=C_{S}\big{(}D(G),f;A\big{)},\quad\quad P\big{(}D(G),f;A\big{)}=P_{\emptyset}\big{(}D(G),f;A\big{)}$
and
$C_{\emptyset}\big{(}D(G),f;A\big{)}=\sum_{S\subseteq E(G)}P_{S}\big{(}D(G),f;A\big{)}.$
It follows from (2.2) that
$P\big{(}D(G),f;A\big{)}=\sum_{S\subseteq E(G)}(-1)^{|S|}C_{S}\big{(}D(G),f;A\big{)}=\sum_{S\subseteq E(G)}(-1)^{|S|}C_{S}\big{(}D(G)|S,f;A\big{)}.$ (2.3)
In addition, [11, Proposition 2.2] demonstrates that $G$ admits a vertex coloring $c$ such that $c(v)-c(u)=f(e)$ for any $e=uv\in E(G)$ with $D(e)=(u,v)$ if and only if the function $f$ satisfies the condition $\sum_{e\in E(C)}\eta_{D(C)}(e)f(e)=0$ for every cycle $C$ of $G$. This condition is equivalent to $\alpha(C)=0$ for all cycles $C$ of $G$ since $\alpha=\alpha_{D,f}$. Therefore, $C_{S}\big{(}D(G)|S,f;A\big{)}=0$ whenever $G|S$ is an $\alpha$-incompatible subgraph of $G$. Then, $P\big{(}D(G),f;A\big{)}$ in (2.3) can be simplified to the following form
$P\big{(}D(G),f;A\big{)}=\sum_{\begin{subarray}{c}S\subseteq E(G),\\ G|S\mbox{ is $\alpha_{G|S}$-compatible}\end{subarray}}(-1)^{|S|}C_{S}\big{(}D(G)|S,f;A\big{)}.$ (2.4)
Denote by $C_{S}\big{(}D(G)|S,f_{S};A\big{)}$ the collection of vertex colorings $c:V(G)\to A$ of $G|S$ such that $c(v)-c(u)=f(e)$ for each edge $e=uv\in S$ with $D(e)=(u,v)$. Then, for a fixed vertex coloring $c\in C_{S}\big{(}D(G)|S,f_{S};A\big{)}$, every vertex coloring $c^{\prime}\in C_{S}\big{(}D(G)|S,f_{S};A\big{)}$ can gives rise to a unique vertex coloring $c^{\prime}-c\in C_{S}\big{(}D(G)|S,0_{S};A\big{)}$. Conversely, every vertex coloring $c^{\prime\prime}\in C_{S}\big{(}D(G)|S,0_{S};A\big{)}$ also gives rise to a unique vertex coloring $c+c^{\prime\prime}\in C_{S}\big{(}D(G)|S,f_{S};A\big{)}$. Notice
$C_{S}\big{(}D(G)|S,f;A\big{)}=|C_{S}\big{(}D(G)|S,f_{S};A\big{)}|\quad\mbox{\rm~{}and~{}}\quad C_{S}\big{(}D(G)|S,0;A\big{)}=|C_{S}\big{(}D(G)|S,0_{S};A\big{)}|.$
It follows that
$C_{S}\big{(}D(G)|S,f;A\big{)}=C_{S}\big{(}D(G)|S,0;A\big{)}=|A|^{c(G|S)}.$
Together with (2.4), we have
$P\big{(}D(G),f;A\big{)}=\sum_{\begin{subarray}{c}S\subseteq E(G),\\ G|S\mbox{ is $\alpha_{G|S}$-compatible}\end{subarray}}(-1)^{|S|}|A|^{c(G|S)},$
which completes the proof. ∎
As a direct consequence of Theorem 2.1, the cycle-assigning polynomials corresponding to admissible assignings precisely count the number of $(A,f)$-colorings of $G$ for any Abelian group $A$ and function $f:E(G)\to A$.
Theorem 2.2.
Let $\alpha$ be an admissible assigning of $G$. Then $P(G,\alpha;k)=P\big{(}D(G),f;A\big{)}$ for every orientation $D$ of $G$, every Abelian group $A$ of order $k$ and every function $f:E(G)\to A$ satisfying $\alpha_{D,f}=\alpha$.
By applying Theorem 2.2, we obtain the next corollary, which states that the number of $(G,f)$-colorings depends only on $\alpha_{D,f}$, and not on the structure of $A$ and the choice of the orientation $D$.
[…]
Let $D,D^{\prime}$ be distinct orientations of $G$, and $A,A^{\prime}$ be two Abelian groups of the same order and functions $f:E(G)\to A$, $f^{\prime}:E(G)\to A^{\prime}$. If $\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$, then $P\big{(}D(G),f;A\big{)}=P\big{(}D^{\prime}(G),f^{\prime};A^{\prime}\big{)}$.
Proof.
By Theorem 2.2, we have
$P\big{(}D(G),f;A\big{)}=P(G,\alpha;k)=P\big{(}D^{\prime}(G),f^{\prime};A^{\prime}\big{)},$
where $\alpha=\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$. ∎
We further explore the fundamental properties of cycle-assigning polynomials. Suppose $e\in E(G)$ is a link of $G$. Then the set $\mathcal{C}(G/e)$ of cycles of $G/e$ consists of the following two parts:
$\mathcal{C}_{-e}(G):=\big{\{}C\in\mathcal{C}(G/e):C\in\mathcal{C}(G),e\notin E(C)\big{\}},\;\mathcal{C}_{/e}(G/e):=\big{\{}C/e:C\in\mathcal{C}(G),e\in E(C)\big{\}}.$
In this sense, each assigning $\alpha$ of $G$ induces an assigning $\alpha_{G/e}$ of $G/e$, where $\alpha_{G/e}(C):=\alpha(C)$ for $C\in\mathcal{C}_{-e}(G)$ and $\alpha_{G/e}(C/e):=\alpha(C)$ for $C\in\mathcal{C}(G)$ containing $e$. We first obtain a deletion-contraction formula for general cycle-assigning polynomials.
Theorem 2.3 (Deletion-contraction Formula).
Let $\alpha$ be an assigning of $G$. Then the cycle-assigning polynomial satisfies
$P(G,\alpha;k)=\begin{cases}k^{|V(G)|},&\mbox{ if }E(G)=\emptyset;\\ \frac{k-1}{k}P(G-e,\alpha_{G-e};k),&\mbox{ if }e\mbox{ is a bridge};\\ \alpha(e)P(G-e,\alpha_{G-e};k),&\mbox{ if }e\mbox{ is a loop};\\ P(G-e,\alpha_{G-e};k)-P(G/e,\alpha_{G/e};k),&\mbox{ otherwise}.\end{cases}$
Proof.
Let $\mathcal{S}_{e}$ be the family of $\alpha$-compatible spanning subgraphs of $G$ containing $e$ and $\mathcal{S}_{-e}$ be the family of $\alpha$-compatible spanning subgraphs of $G$ that do not contain $e$. If $e$ is a bridge, then a subgraph $H$ containing $e$ is an $\alpha$-compatible spanning subgraph of $G$ if and only if $H-e$ is an $\alpha$-compatible spanning subgraph of $G$. Therefore, we conclude that every $\alpha$-compatible spanning subgraph $H$ of $G$ containing $e$ gives rise to a unique $\alpha$-compatible spanning subgraph $H-e$ of $G$ that does not contain $e$, and vice versa. Then we have
$P(G,\alpha;k)=\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H+e)|}k^{c(H+e)}+\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H)|}k^{c(H)}.$
Since $e$ is a bridge, $c(H+e)=c(H)-1$ for any $H\in\mathcal{S}_{-e}$. Hence, we deduce
$P(G,\alpha;k)=\frac{k-1}{k}P(G-e,\alpha_{G-e};k).$
If $e$ is a loop, we need to consider both cases: $\alpha(e)=0$ and $\alpha(e)=1$. For the former case, similarly to the case where $e$ is a bridge, we can easily derive $P(G,\alpha;k)=0$. For the latter case, it is clear that $H$ is an $\alpha$-compatible spanning subgraph of $G$ if and only if $H$ is an $\alpha_{G-e}$-compatible spanning subgraph of $G-e$. So, we arrive at $P(G,\alpha;k)=P(G-e,\alpha_{G-e};k)$.
If $e$ is neither a bridge nor a loop, then $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing no $e$ if and only if $H$ is an $\alpha_{G-e}$-compatible spanning subgraph of $G-e$, and $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing $e$ if and only if $H/e$ is an $\alpha_{G/e}$-compatible spanning subgraph of $G/e$. Therefore, we have
$P(G,\alpha;k)$ $=\sum_{H\in\mathcal{S}_{-e}}(-1)^{|E(H)|}k^{c(H)}+\sum_{H\in\mathcal{S}_{e}}(-1)^{|E(H)|}k^{c(H)}$
$=P(G-e,\alpha_{G-e};k)+\sum_{H\in\mathcal{S}_{e}}(-1)^{|E(H/e)|+1}k^{c(H/e)}$
$=P(G-e,\alpha_{G-e};k)-P(G/e,\alpha_{G/e};k),$
which finishes the proof. ∎
In addition, the cycle-assigning polynomial $P(G,\alpha;k)$ has the following simple decomposition formula.
Theorem 2.4.
Let $\alpha$ be an assigning of $G$ and $G=G_{1}\sqcup G_{2}\sqcup\cdots\sqcup G_{i}$, where $G_{j}$ is the component of $G$ for $j=1,2,\ldots,i$. Then
$P(G,\alpha;k)=\prod_{j=1}^{i}P(G_{j},\alpha_{G_{j}};k).$
Proof.
Since $G$ decomposes into the union of the components $G_{j}$ of $G$, each $\alpha$-compatible spanning subgraph of $G$ is partitioned into the disjoint union of the $\alpha_{G_{j}}$-compatible spanning subgraph of $G_{j}$, and vice versa. Therefore, we can deduce $P(G,\alpha;k)=\prod_{j=1}^{i}P(G_{j},\alpha_{G_{j}};k)$, which finishes the proof. ∎
At the end of the section, we list the cycle-assigning polynomials for the empty graph, a tree and the cycle on $n$ vertices.
[…]
  • (1)
    Let $O_{n}$ be the empty graph on $n$ vertices. Then $\alpha\equiv 0$ and
    $P(O_{n},0;k)=k^{n}.$
  • (2)
    Let $T_{n}$ be a tree on $n$ vertices. Then $\alpha\equiv 0$ and
    $P(T_{n},0;k)=k(k-1)^{n-1}=\sum_{i=1}^{n}(-1)^{n-i}\binom{n-1}{i-1}k^{i}.$
  • (3)
    Let $C_{n}$ be a cycle on $n$ vertices. Combining part (2) and Theorem 2.3, we have
    $P(C_{n},\alpha;k)=\begin{cases}(k-1)^{n}+(-1)^{n}(k-1),&\mbox{ if }\alpha(C_{n})=0;\\ (k-1)^{n}+(-1)^{n+1},&\mbox{ if }\alpha(C_{n})=1.\end{cases}$
3 Polynomials counting nowhere-zero chains associated with homomorphisms
In this section, we mainly clarify the close relation between the cycle-assigning polynomial and the $\alpha$-assigning polynomial, which was first introduced by Kochol [13] to count nowhere-zero chains in graphs–nonhomogeneous analogues of nowhere-zero flows. Most recently, Kochol [14] further extended the approach to regular matroids. For this purpose, we first review some basic definitions about regular chain groups associated to regular matroids. For additional properties of regular chain groups, refer to the references [1, 29].
A regular matroid $M=M(U)$ on a ground set $E$ is represented by a totally unimodular matrix $U$. A more general setting and background on matroids can be found in Oxley’s book [25]. A chain on $E$ over an Abelian group $A$ is a map $f\in A^{E}$, and its support $\sigma(f)$ is $\sigma(f):=\{f(e)\neq 0\mid e\in E\}$. We say that $f$ is proper if $\sigma(f)=E$. Associated with $U$ or $M(U)$, a regular chain group $N$ on $E$ over $\mathbb{Z}$ is the set of integral vectors orthogonal to each row of $U$. The set of chains orthogonal to every chain of $N$ forms a chain group called orthogonal to $N$, denoted by $N^{\perp}$. Actually, $N^{\perp}$ consists of the integral vectors from the line combinations of the rows of $U$. A chain $f$ of $N$ is elementary if there is a no nonzero chain $g$ of $N$ such that $\sigma(g)\subset\sigma(f)$. An elementary chain $f$ of $N$ is said to be primitive if $f(e)\in\{\pm 1,0\}$ for all $e\in E$. In this section, we always assume that a regular chain group $N$ is associated with a totally unimodular matrix $U(N)$ representing a regular matroid $M=M(N)$.
Let $A$ be an additive Abelian group. We shall consider $A$ as a (right) $\mathbb{Z}$-module such that the scalar multiplication $a\cdot z$ of $a\in A$ by $z\in\mathbb{Z}$ equals zero if $z=0$, $\sum_{1}^{z}a$ if $z>0$, and $\sum_{1}^{-z}(-a)$ if $z<0$. Analogously, if $a\in A$ and $f\in\mathbb{Z}^{E}$, then define $a\cdot f\in A^{E}$ such that $(a\cdot f)(e)=a\cdot f(e)$ for every $e\in E$. We shall denote by $A(N)$ the $A$-chain group generated by the regular chain group $N$, that is,
$A(N):=\Big{\{}\sum_{i=1}^{m}a_{i}\cdot f_{i}\mid a_{i}\in A,f_{i}\in N,m\geq 1\Big{\}}.$
Let $A[N]$ be the collection of all proper $A$-chains from $A(N)$, i.e.,
$A[N]:=\big{\{}f\in A(N)\mid\sigma(f)=E\big{\}}.$
From [1, Proposition 1], we have
$g\in A^{E}\text{ is from }A(N)\Longleftrightarrow\sum_{e\in E}g(e)\cdot f(e)=0\text{ for any }f\in N^{\perp}.$ (3.1)
Let $\psi$ be a homomorphism from a regular chain group $N$ to $A$. Define by
$A_{\psi}[N]:=\Big{\{}g\in(A-0)^{E}\mid\sum_{e\in E}g(e)\cdot f(e)=\psi(f)\text{ for all }f\in N\Big{\}}.$
Each element from $A_{\psi}[N]$ is called a proper $(N,\psi)$-chain.
Let $\mathcal{P}(N)$ be the set of primitive chains of a regular chain group $N$, and $\mathcal{C}(M)$ be the family of circuits of $M=M(N)$. By [29], $\mathcal{P}(N)$ and $\mathcal{C}(M)$ have the following relationship
$\mathcal{C}(M)=\big{\{}\sigma(c)\mid c\in\mathcal{P}(N)\big{\}}.$ (3.2)
An assigning of $M$ is any mapping $\alpha$ from $\mathcal{C}(M)$ to $\{0,1\}$. Write $\alpha\equiv 0$ if $\alpha(C)=0$ for all $C\in\mathcal{C}(M)$ or $\mathcal{C}(M)=\emptyset$. Let $e\in E$. It follows from the relation $\mathcal{C}(M-e)\subseteq\mathcal{C}(M)$ that an assigning $\alpha$ of $M$ naturally gives rise to an assigning $\alpha[-e]$ of $M-e$ for which
$\alpha[-e](C)=\alpha(C),\quad C\in\mathcal{C}(M-e).$
By [25, Proposition 3.1.10], we know that
if $C\in\mathcal{C}(M/e)$, then either $C\in\mathcal{C}(M)$ or $C\cup e\in\mathcal{C}(M)$.
Similarly, the assigning $\alpha$ of $M$ also induces an assigning $\alpha[/e]$ of $M/e$ such that $\alpha[/e](C)=\alpha(C)$ if $C\in\mathcal{C}(M/e)\cap\mathcal{C}(M)$, and $\alpha[/e](C)=\alpha(C\cup e)$ otherwise.
Moreover, from (3.2), the homomorphism $\psi$ automatically leads to an assigning $\alpha_{N,\psi}$ of $M$ such that for every $c\in\mathcal{P}(N)$,
$\alpha_{N,\psi}\big{(}\sigma(c)\big{)}:=\begin{cases}0,&\text{if }\psi(c)=0;\\ 1,&\text{if }\psi(c)\neq 0.\end{cases}$
An assigning $\alpha$ of $M$ is said to be homogeneous if there is a homomorphism $\psi:N\to A$ such that $\alpha=\alpha_{N,\psi}$. Furthermore, $\psi$ is proper and $\alpha=\alpha_{N,\psi}$ is called a proper assigning of $M$ if $A_{\psi}[N]\neq\emptyset$.
In 2024, Kochol showed in [13, Theorem 1] that there exists a polynomial $p(M,\alpha;k)$ counting the number of nowhere-zero chains in $A_{\psi}[N]$. The polynomial $p(M,\alpha;k)$ is called an $\alpha$-assigning polynomial of $M$ by Kochol.
Theorem 3.1 ([14], Theorem 1).
Suppose $M$ is a regular matroid on $E$ and $\alpha$ is a homogeneous assigning of $M$. Then, there exists a polynomial $p(M,\alpha;k)$, such that $p(M,\alpha;k)=|A_{\psi}[N]|$ for every regular chain group $N$ on $E$ and $M(N)=M$, every Abelian group $A$ of order $k$, and every homomorphism $\psi:N\to A$ satisfies $\alpha_{N,\psi}=\alpha$. If $\alpha$ is proper, then $p(M,\alpha;k)$ has degree $r(M)$, and $p(M,\alpha;k)=0$ otherwise. Furthermore, for any $e\in E$, $\alpha[-e]$ and $\alpha[/e]$ are homogeneous assignings of $M-e$ and $M/e$, respectively, and
$p(M,\alpha;k)=\begin{cases}1,&\text{if $E=\emptyset$};\\ (k-1)p\big{(}M-e,\alpha[-e];k\big{)},&\text{if $e$ is a bridge};\\ \alpha(e)p\big{(}M-e,\alpha[-e];k\big{)},&\text{if $e$ is a loop};\\ p\big{(}M-e,\alpha[-e];k\big{)}-p\big{(}M/e,\alpha[/e];k\big{)},&\text{otherwise}.\\ \end{cases}$
Below we shall explore some applications of the $\alpha$-assigning polynomial to colorings and tensions of $G$. The incidence matrix of $D(G)$ is the $|V(D)|\times|A(D)|$ matrix $M_{D(G)}:=(m_{va})$ whose rows and columns are indexed by the vertices and arcs, where, for a vertex $v$ and arc $a$
$m_{va}:=\begin{cases}1,&\text{ if arc }a\text{ is a link and vertex }v\text{ is the head of }a;\\ -1,&\text{ if arc }a\text{ is a link and vertex }v\text{ is the tail of }a;\\ 0,&\text{ otherwise}.\end{cases}$
If $c\in A^{V(G)}$, then the coboundary of $c$, with respect to the orientation $D$, is the map $\delta c\in A^{E(G)}$ given by
$\delta c(e):=\sum_{v\in V(G)}m_{vD(e)}c(v).$
In this sense, we call a mapping $t\in A^{E(G)}$ a tension or an $A$-tension if there exists a vertex coloring $c\in A^{V(G)}$ such that $t=\delta c$. Similar to $(A,f)$-coloring associated with a mapping $f\in A^{E(G)}$, a tension $t$ is referred to as an $(A,f)$-tension if it satisfies $t(e)\neq f(e)$ for all $e\in E(G)$. In particular, taking $f\equiv 0$, the $(A,0)$-tension is exactly the nowhere-zero $A$-tension. It is clear that all $A$-tensions form an (additive) Abelian subgroup of $A^{E(G)}$, denoted by $\Delta(G,A)$. Let $\Delta_{f}(G,A)$ be the set of all $(A,f)$-tensions of $G$ for $f\in A^{E(G)}$. The number of nowhere-zero tensions is evaluated by a polynomial $\tau(G;k)$ in $k$ introduced by Tutte [28]. This was developed and formally named the tension polynomial of $G$ by Kochol [11] in 2002. Motivated by their work, a natural question arises:
  • Is there a polynomial function of $|A|$ that generalizes the tension polynomial and counts $(A,f)$-tensions in graphs?
The answer to the question was implicit in the work of Kochol [14]. By applying Theorem 3.1, we shall present a precise answer in Section 3.
Let $M(G)$ be the cycle matroid of $G$ represented by the incidence matrix $M_{D(G)}$ ($M(G)$ is a regular matroid, see [25, Proposition 5.1.5]), and $\Gamma(G)$ be the regular chain group on $E(G)$ associated with $M_{D(G)}$. From [29], the mapping $\eta_{D(C)}$ defined in (2.1) is the primitive chain of $\Gamma(G)$ for every cycle $C$ of $G$, and $\Gamma(G)$ is generated by $\big{\{}\eta_{D(C)}\mid C\in\mathcal{C}(G)\big{\}}$. In addition, the regular chain group $\Delta(G)$ on $E(G)$ over $\mathbb{Z}$ consisting of all line combinations of rows from the matrix $M_{D(G)}$ is exactly orthogonal to $\Gamma(G)$, i.e., $\Delta(G)=\Gamma(G)^{\perp}$. Again, from [29], the chain group $\Delta(G,A)$ can be generated by $\Delta(G)$. Then we have the following result.
[…]
Let $\alpha$ be an admissible assigning of $G$. Then $\tau\big{(}G,\alpha;k\big{)}$ and $p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}$ are the same polynomial, and they satisfy that
$\tau\big{(}G,\alpha;k\big{)}=p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}=|A_{\psi}[\Gamma(G)]|=|\Delta_{f}(G,A)|$
for every Abelian group $A$ of order $k$, every function $f:E(G)\to A$ satisfying $\alpha_{D,f}=\alpha$, and every homomorphism $\psi:\Gamma(G)\to A$ such that $\psi\big{(}\eta_{D(C)}\big{)}=-\sum_{e\in E(G)}f(e)\cdot\eta_{D(C)}(e)$ for any cycle $C$ of $G$. Furthermore, $\tau\big{(}G,\alpha;k\big{)}$ has degree $|V(G)|-c(G)$ if $\alpha(C)=1$ for any loop $C$ of $G$ and equals zero otherwise, and $\tau\big{(}G,\alpha;k\big{)}$ also satisfies
$\tau(G,\alpha;k)=\begin{cases}1,&\text{if $E(G)=\emptyset$};\\ (k-1)\tau(G-e,\alpha_{G-e};k),&\text{if $e$ is a bridge};\\ \alpha(e)\tau(G-e,\alpha_{G-e};k),&\text{if $e$ is a loop};\\ \tau\big{(}G-e,\alpha_{G-e};k\big{)}-\tau(G/e,\alpha_{G/e};k),&\text{otherwise}.\\ \end{cases}$
To prove Section 3, the following lemmas are required.
[…]
Given a mapping $f\in A^{E(G)}$. Let
$A_{f}[G]=\big{\{}g\in(A-0)^{E(G)}\mid g+f\in\Delta(G,A)\big{\}}.$
Then the mapping $\phi:\Delta_{f}(G,A)\to A_{f}[G]$ sending $g\in\Delta_{f}(G,A)$ to $\phi(g)=g-f$ is a bijection.
Proof.
Taking any element $g\in\Delta_{f}(G,A)$, by $g(e)\neq f(e)$ for all $e\in E(G)$, then $\phi(g)\in(A-0)^{E(G)}$ and $\phi(g)+f=g\in\Delta_{f}(G,A)\subseteq\Delta(G,A)$. Namely, $\phi$ is well-defined. Clearly, $\psi$ is injective.
Conversely, define the mapping $\phi^{-1}:A_{f}[G]\to\Delta_{f}(G,A)$ sending $g\in A_{f}[G]$ to $\phi^{-1}(g)=g+f$. Given an element $g\in A_{f}[G]$. Since $g(e)\neq 0$ for every $e\in E(G)$, $\phi^{-1}(g)\in\Delta_{f}(G,A)$. Thus, $\phi^{-1}$ is well-defined. The mapping $\phi^{-1}$ is an injection obviously. Therefore, $\phi$ is bijective. We complete the proof. ∎
[…]
Suppose $f\in A^{E(G)}$ and the homomorphism $\psi:\Gamma(G)\to A$ such that $\psi\big{(}\eta_{D(C)}\big{)}=-\sum_{e\in E(G)}f(e)\cdot\eta_{D(C)}(e)$ for any cycle $C$ of $G$. Then $A_{\psi}[\Gamma(G)]=A_{f}[G]$.
Proof.
We first show $A_{\psi}[\Gamma(G)]\subseteq A_{f}[G]$. Note that the chain group $\Delta(G,A)$ is generated by the regular chain group $\Delta(G)$ and $\Delta(G)$ is orthogonal to $\Gamma(G)$. Then, from (3.1), this proof reduces to verifying that for any $g\in A_{\psi}[\Gamma(G)]$, $\sum_{e\in E(G)}(g+f)(e)\cdot h(e)=0$ for all $h\in\Gamma(G)$. Since $\Gamma(G)$ is generated by $\big{\{}\eta_{D(C)}\mid C\in\mathcal{C}(G)\big{\}}$, the proof is further equivalent to checking that for any $g\in A_{\psi}[\Gamma(G)]$,
$\sum_{e\in E(G)}(g+f)(e)\cdot\eta_{D(C)}(e)=0,\quad\forall\,C\in\mathcal{C}(G).$
Since $\sum_{e\in E(G)}g(e)\cdot\eta_{D(C)}(e)=\psi\big{(}\eta_{D(C)}\big{)}$ for every $g\in A_{\psi}[\Gamma(G)]$, it follows from $\psi\big{(}\eta_{D(C)}\big{)}=-\sum_{e\in E(G)}f(e)\cdot\eta_{D(C)}(e)$ that $\sum_{e\in E(G)}(g+f)(e)\cdot\eta_{D(C)}(e)$ is indeed equal to zero. This implies that $g+f\in\Delta(G,A)$ for any $g\in A_{\psi}[\Gamma(G)]$, that is, $A_{\psi}[\Gamma(G)]\subseteq A_{f}[G]$.
For the opposite inclusion, for any $g\in A_{f}[G]$, we have $g+f\in\Delta(G,A)$. Again, using (3.1), we have $\sum_{e\in E(G)}(g+f)(e)\cdot\eta_{D(C)}(e)=0$ for all cycles $C$ of $G$. It follows that
$\sum_{e\in E(G)}g(e)\cdot\eta_{D(C)}(e)=-\sum_{e\in E(G)}f(e)\cdot\eta_{D(C)}(e)=\psi\big{(}\eta_{D(C)}\big{)}.$
This means $g\in A_{\psi}[\Gamma(G)]$ for each $g\in A_{f}[G]$. Namely, we have $A_{\psi}[\Gamma(G)]\supseteq A_{f}[G]$. Then we arrive at $A_{\psi}[\Gamma(G)]=A_{f}[G]$, which finishes the proof. ∎
Now we proceed to prove Section 3.
Proof of Section 3.
Note that any $(A,f)$-coloring $c$ of $G$ yields an $(A,f)$-tension $\delta c$ of $G$, and vice versa. We assert that for every $(A,f)$-tension $t$, there are exactly $|A|^{c(G)}$ $(A,f)$-colorings $c$ satisfying $\delta c=t$. Suppose $t$ is an $(A,f)$-tension of $G$ satisfying $t=\delta c$ for some $c\in A^{V(G)}$. We construct a mapping $c^{\prime}\in A^{V(G)}$ such that $c^{\prime}(u)-c(u)=c^{\prime}(v)-c(v)$ whenever vertices $u$ and $v$ belong to the same component of $G$. It follows that $c^{\prime}$ is a $(A,f)$-coloring and $\delta c^{\prime}=t$. On the other hand, if $\delta c^{{}^{\prime\prime}}=t$ for another $(A,f)$-coloring $c^{{}^{\prime\prime}}$, then $c^{{}^{\prime\prime}}(u)-c(u)=c^{{}^{\prime\prime}}(v)-c(v)$ whenever $u$ and $v$ are in the same component of $G$. Thus, we have shown the assertion. It follows from Theorem 2.2 that
$\tau\big{(}G,\alpha;k\big{)}=|\Delta_{f}(G,A)|.$ (3.3)
Applying Theorem 3.1, we have
$p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}=|A_{\psi}[\Gamma(G)]|.$ (3.4)
In addition, by Section 3 and Section 3, we obtain
$|\Delta_{f}(G,A)|=|A_{\psi}[\Gamma(G)]|.$ (3.5)
Immediately, by integration of (3.3), (3.4) and (3.5), we arrive at
$\tau\big{(}G,\alpha;k\big{)}=p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}=|A_{\psi}[\Gamma(G)]|=|\Delta_{f}(G,A)|.$
It should be noted that both $\tau(G,\alpha;k)$ and $p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}$ are also polynomials in $k$ of degree at most $|V(G)|-c(G)$ and these two polynomials agree on infinitely many positive integers $k$. Hence, from the fundamental theorem of algebra, $\tau(G,\alpha;k)$ and $p\big{(}M(G),\alpha_{\Gamma(G),\psi};k\big{)}$ are the same polynomial.
In fact, we have $\alpha_{\Gamma(G),\psi}=\alpha_{D,f}=\alpha$ from (2.1). Moreover, [14, Theorem 3] states that
a homogeneous assigning α of M(G) is proper  α(e)=1 for each loop e of $\text{a homogeneous assigning $\alpha$ of $M(G)$ is proper $\Longleftrightarrow$ $\alpha(e)=1$ for each loop $e$ of $M(G)$}.$
Together with Section 3 and $r(M(G))=|V(G)|-c(G)$, we have that $\tau\big{(}G,\alpha;k\big{)}$ has degree $|V(G)|-c(G)$ if $\alpha(C)=1$ for any loop $C$ of $G$ and equals zero otherwise. Finally, the remaining part is a direct result of Theorem 2.3 via $\tau(G,\alpha;k)=k^{-c(G)}P(G,\alpha;k)$. We complete the proof. ∎
Notably, [11, Proposition 2.2] states that $f\in A^{E(G)}$ is a tension of $G$ if and only if $f$ satisfies the condition $\sum_{e\in E(C)}\eta_{D(C)}(e)f(e)=0$ for every cycle $C$ of $G$. It follows that $\alpha_{D,f}\equiv 0$ for every $f\in\Delta(G,A)$. So, the assigning polynomial $\tau(G,\alpha_{D,f};k)$ is the classical tension polynomial of $G$ (see [11]) whenever $f\in\Delta(G,A)$.
Let $M$ be a matroid on the ground set $E$ and $\prec$ be a linear ordering of $E$. For every subset $X\subseteq E$, the notation ${\rm min}(X)$ represents the minimal element of $X$ with respect to $\prec$. A subset $X\subseteq E$ is $(M,\prec)$-compatible if $C\cap X\neq\{{\rm min}(C)\}$ for each $C\in\mathcal{C}(M)$, where $\mathcal{C}(M)$ is the set of all circuits of $M$. Furthermore, we use $\mathcal{E}(M,\prec)$ to denote the collection of all $(M^{*},\prec)$-compatible subsets of $E$, where $M^{*}$ is the dual matroid of $M$. For any $X\subseteq E$ and assigning $\alpha$ of $M$, we define $\delta(M,\alpha;X)$ to be
$\delta(M,\alpha;X)=:\begin{cases}0,&\text{if there exists $C\in\mathcal{C}(M)$such that $C\subseteq X$ and $\alpha(C)=1$};\\ 1,&\text{otherwise}.\end{cases}$
For a regular matroid $M$, Kochol derived an explicit expression for the $\alpha$-assigning polynomial $p(M,\alpha;k)$ over all $(M^{*},\prec)$-compatible subsets of $E$ in [14, Theorem 2].
Theorem 3.2 ([14], Theorem 2).
Let $\alpha$ be a homogeneous assigning of a regular matroid $M$ on $E$ and $\prec$ be a linear ordering of $E$. Then,
$p(M,\alpha;k)=\sum_{X\in\mathcal{E}(M,\prec)}\delta(M,\alpha;X)(-1)^{|X|}(k-1)^{r(M/X)}.$
Below we return to the cycle matroid $M(G)$ on $E(G)$ represented by $M_{D(G)}$. Its dual matroid $M^{*}(G)$, also on $E(G)$, is referred to as a bond matroid of $G$ ($G$ may be not a planar graph). [25, Proposition 2.3.1] demonstrates that
[…]
A bond of a graph is defined as a minimal nonempty edge cut, which is the smallest nonempty subset of edges whose removal from the graph results in an increase in the number of components. Let $\prec$ be a linear ordering of $E(G)$. Define $\mathcal{E}(G,\prec)$ as
$\mathcal{E}(G,\prec):=\big{\{}X\subseteq E(G)\mid X\cap C^{*}\neq\big{\{}{\rm min}(C^{*})\big{\}}\text{ for all bonds $C^{*}$ of $G$}\big{\}}.$
It follows that
$\mathcal{E}(G,\prec)=\mathcal{E}\big{(}M(G),\prec\big{)}.$ (3.6)
For any $X\subseteq E(G)$ and assigning $\alpha$ of $G$, we define $\delta(G,\alpha;X)$ by
$\delta(G,\alpha;X)=:\begin{cases}0,&\text{if there exists $C\in\mathcal{C}(G)$such that $E(C)\subseteq X$ and $\alpha(C)=1$};\\ 1,&\text{otherwise}.\end{cases}$
From (2.1), every admissible assigning $\alpha=\alpha_{D,f}$ of $G$ induced by $f\in A^{E(G)}$ is precisely a homogeneous assigning $\alpha_{\Gamma(G),\psi}$ of $M(G)$, where the homomorphism $\psi:\Gamma(G)\to A$ satisfies $\psi\big{(}\eta_{D(C)}\big{)}=-\sum_{e\in E(G)}f(e)\cdot\eta_{D(C)}(e)$ for any cycle $C$ of $G$. In this sense, we have
$\delta(G,\alpha;X)=\delta\big{(}M(G),\alpha;X\big{)},\quad\text{ for every }X\subseteq E(G).$ (3.7)
By integrating the preceding discussion, as a direct consequence of Section 3 and Theorem 3.2, we conclude an analogous result of Theorem 3.2 concerning $\tau(G,\alpha;k)$ and $P(G,\alpha;k)$ in the following corollary.
[…]
Let $\alpha$ be an admissible assigning of $G$ on $E(G)$ and $\prec$ be a linear ordering of $E(G)$. Then,
$\tau(G,\alpha;k)=\sum_{X\in\mathcal{E}(G,\prec)}\delta(G,\alpha;X)(-1)^{|X|}(k-1)^{r(G)-r(X)}$
and
$P(G,\alpha;k)=\sum_{X\in\mathcal{E}(G,\prec)}\delta(G,\alpha;X)(-1)^{|X|}(k-1)^{c(G|X)},$
where $r$ is the rank function of $G$ given by $r(X):=|V(G)|-c(G|X)$ for $X\subseteq E(G)$.
Proof.
Suppose $\alpha$ is equal to the admissible assigning $\alpha_{D,f}$ for some $f\in A^{E(G)}$. Together with the relations in (3.6) and (3.7), by using Section 3 and Theorem 3.2, we have
$\tau(G,\alpha;k)=\sum_{X\in\mathcal{E}(G,\prec)}\delta(G,\alpha;X)(-1)^{|X|}(k-1)^{r(M(G)/X)}.$
Note that $r\big{(}M(G)/X\big{)}=r\big{(}M(G)\big{)}-r_{M(G)}(X)$ from [25, Proposition 3.1.6], where $r_{M(G)}$ is the rank function of $M(G)$. It follows from $r_{M(G)}(X)=r(X)$ for any $X\subseteq E(G)$ that
$\tau(G,\alpha;k)=\sum_{X\in\mathcal{E}(G,\prec)}\delta(G,\alpha;X)(-1)^{|X|}(k-1)^{r(G)-r(X)}.$
Combining the relations $r(X)=|V(G)|-c(G|X)$ and $P(G,\alpha;k)=k^{c(G)}\tau(G,\alpha;k)$, we derive the second formula in Section 3. We finish the proof. ∎
4 Generalization of Whitney’s Broken Cycle Theorem
A natural question is whether there is a combinatorial interpretation for the coefficients of cycle-assigning polynomials. To address this, we introduce the concept of a broken $\alpha$-compatible cycle as an extension of the usual broken cycle proposed by Whitney [30].
[…]
Let $\alpha$ be an assigning of $G$ and $\prec$ be a linear ordering on $E(G)$. We say that a subgraph $H$ of $G$ is said to be:
  • an $\alpha$-compatible cycle of $G$ if $H$ is a cycle and $\alpha(H)=0$;
  • a broken $\alpha$-compatible cycle of $G$ if $H$ is obtained from an $\alpha$-compatible cycle by removing the maximal element with respect to the linear ordering $\prec$.
[…]
Let $\alpha$ be a $D$-admissible assigning of $G$ and $\prec$ a linear ordering on $E(G)$. Suppose $H$ is a subgraph of $G$ containing a broken $\alpha$-compatible cycle $B$ and $B=C-e_{C}$ for some $\alpha$-compatible cycle $C$ of $G$, where $e_{C}$ is the maximal member in $E(C)$ with respect to $\prec$. If $H$ is an $\alpha$-compatible subgraph of $G$ and $e_{C}\notin E(H)$, then $H+e_{C}$ is also an $\alpha$-compatible subgraph of $G$.
Proof.
Note that $C$ is a cycle in $H+e_{C}$, but not in $H$. If $H+e_{C}$ contains exactly one more cycle $C$ than $H$, there is nothing left to prove. Otherwise, $H+e_{C}$ contains at least one additional cycle compared to $H$, excluding the cycle $C$. Notably, every such additional cycle, together with the cycle $C$, forms a theta-subgraph in $H+e_{C}$. Therefore, this case reduces to verifying that each such the theta-subgraph in $H+e_{C}$ is an $\alpha$-compatible subgraph of $G$. Notice that $\alpha$ is a $D$-admissible assigning of $G$. Without loss of generality, we only need to examine the following oriented theta-subgraph in $H+e_{C}$ with the orientation $D$, as shown in Figure 1.
[…]
Figure 1: The directed paths $P_{1},P_{2}$ and $P_{3}$ toghter form an oriented theta-graph
More specifically, in Figure 1, the edge $e_{C}=u_{0}v_{0}$; the cycle $C$, which is the union of the red directed path $P_{1}=(u_{i},u_{i-1},\ldots,v_{k-1},v_{k})$ and the green directed path $P_{2}=(v_{k},v_{k+1},\ldots,v_{l},u_{i})$, and the cycle $C_{1}$, which is the union of the red directed path $P_{1}$ and the blue directed path $P_{3}=(u_{i},u_{i+1},\ldots,u_{j},v_{k})$, are in $H+e_{C}$ but not in $H$; the cycle $C_{2}$ is the common cycle of $H+e_{C}$ and $H$, formed by the union of both directed paths $P_{2}$ and $P_{3}$. Then, to complete our proof, it suffices to show $\alpha(C)=\alpha(C_{1})=\alpha(C_{2})=0$.
Since $\alpha$ is $D$-admissible, we may assume $\alpha=\alpha_{D,f}$ for some Abelian group $A$ and function $f:E(G)\to A$. Combining that $C$ and $H$ are $\alpha$-compatible subgraphs of $G$, this means
[…]
This is further equivalent to satisfying the following equations
$\sum_{e\in E(C)}\eta_{D(C)}(e)f(e)=\sum_{e\in E(P_{1})}\eta_{D(C)}(e)f(e)+\sum_{e\in E(P_{2})}\eta_{D(C)}(e)f(e)=0$
and
$\sum_{e\in E(C_{2})}\eta_{D(C_{2})}(e)f(e)=\sum_{e\in E(P_{2})}\eta_{D(C_{2})}(e)f(e)+\sum_{e\in E(P_{3})}\eta_{D(C_{2})}(e)f(e)=0.$
Together with $\eta_{D(C_{1})}(e)=-\eta_{D(C)}(e)$ for any edge $e\in E(P_{1})$ and $\eta_{D(C_{1})}(e)=\eta_{D(C_{2})}(e)$ for any edge $e\in E(P_{3})$, we deduce
$\sum_{e\in E(C_{1})}\eta_{D(C_{1})}(e)f(e)=\sum_{e\in E(P_{1})}\eta_{D(C_{1})}(e)f(e)+\sum_{e\in E(P_{3})}\eta_{D(C_{1})}(e)f(e)=0.$
This implies $\alpha(C_{1})=\alpha_{D,f}(C_{1})=0$. Therefore, we conclude that $H+e_{C}$ is an $\alpha$-compatible subgraph of $G$, which finishes the proof. ∎
The upcoming theorem expresses the coefficients of $P(G,\alpha;k)$ in terms of $\alpha$-compatible spanning subgraphs that do not contain broken $\alpha$-compatible cycles, which is an extended version of the Whitney’s celebrated Broken Cycle Theorem given by Whitney [30]. It is worth noting that the coefficients $w_{i}(G,\alpha)$ in Theorem 4.1 are independent of the choice of edge order.
Theorem 4.1 (Generalized Whitney’s Broken Cycle Theorem).
Let $\alpha$ be an admissible assigning of $G$ and $\prec$ a linear ordering on $E(G)$. Write $P(G,\alpha;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G,\alpha)k^{|V(G)|-i}$. Then each unsigned coefficient $w_{i}(G,\alpha)$ is the number of $\alpha$-compatible spanning subgraphs of $G$ that have exactly $i$ edges and do not contain broken $\alpha$-compatible cycles with respect to $\prec$.
Proof.
Suppose $G$ has $q\geq 0$ broken $\alpha$-compatible cycles. For every broken $\alpha$-compatible cycle $B$, let $e_{B}$ be the maximal member in $B$ with respect to the linear ordering $\prec$. We can arrange the broken $\alpha$-compatible cycles $B_{1},\ldots,B_{q}$ such that $e_{B_{1}}\preceq e_{B_{2}}\cdots\preceq e_{B_{q}}$. Let $\mathcal{S}_{1}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing the broken $\alpha$-compatible cycle $B_{1}$. For $i=2,\ldots,q$, let $\mathcal{S}_{i}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing the broken $\alpha$-compatible cycle $B_{i}$ but not any of $B_{1},\ldots,B_{i-1}$. In addition, let $\mathcal{S}_{q+1}$ be the collection of $\alpha$-compatible spanning subgraphs of $G$ containing no broken $\alpha$-compatible cycles. Then $\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{q+1}\}$ decomposes the set of $\alpha$-compatible spanning subgraphs of $G$. Combining Theorem 2.2, the cycle-assigning polynomial $P(G,\alpha;k)$ can be written in the form
$P(G,\alpha;k)=\sum_{i=1}^{q+1}\sum_{H\in\mathcal{S}_{i}}(-1)^{|E(H)|}k^{c(H)}.$ (4.1)
We first consider $\mathcal{S}_{1}$. Assume $B_{1}=C_{1}-e_{C_{1}}$ for some $\alpha$-compatible cycle $C_{1}$ of $G$, where $e_{C_{1}}$ is the maximal member in $C_{1}$ with respect to $\prec$. Clearly each member $H$ in $\mathcal{S}_{1}$ containing $e_{C_{1}}$ gives rise to a unique member $H-e_{C_{1}}$ in $\mathcal{S}_{1}$ that does not contain $e_{C_{1}}$. Conversely, according to Section 4, each member $H$ in $\mathcal{S}_{1}$ containing no $e_{C_{1}}$ gives rise to a unique member $H+e_{C_{1}}$ in $\mathcal{S}_{1}$ containing $e_{C_{1}}$. In addition, since $B_{1}+e_{C_{1}}=C_{1}$ is a cycle of $G$, for each $H\in\mathcal{S}_{1}$ containing $e_{C_{1}}$, we have $c(H)=c(H-e_{C_{1}})$. Thus, we derive
$\sum_{H\in\mathcal{S}_{1},\,e_{C_{1}}\in E(H)}\big{(}(-1)^{|E(H)|}k^{c(H)}+(-1)^{|E(H)|-1}k^{c(H-e_{C_{1}})}\big{)}=0.$
Namely, we arrive at $\sum_{H\in\mathcal{S}_{1}}(-1)^{|E(H)|}k^{c(H)}=0$ in (4.1).
Consider $\mathcal{S}_{2}$. Assume $B_{2}=C_{2}-e_{C_{2}}$ for some $\alpha$-compatible cycle $C_{2}$ of $G$, where $e_{C_{2}}$ is the maximal member in $C_{2}$ with respect to $\prec$. It follows from the ordering $e_{B_{1}}\preceq e_{B_{2}}$ that $e_{B_{1}}\preceq e_{B_{2}}\prec e_{C_{2}}$. So, $B_{1}$ and $B_{2}$ do not contain $e_{C_{2}}$. Combining this, using the same argument as the case $\mathcal{S}_{1}$, we conclude that every element $H$ in $\mathcal{S}_{2}$ containing no $e_{C_{2}}$ gives rise to a unique member $H+e_{C_{2}}$ in $\mathcal{S}_{2}$ containing $e_{C_{2}}$, and vice versa. Hence, we also obtain $\sum_{H\in\mathcal{S}_{2}}(-1)^{|E(H)|}k^{c(H)}=0$ in (4.1).
Likewise, we can deduce the total contribution of the members in $\mathcal{S}_{i}$ to $P(G,\alpha;k)$ is zero for each $i=1,2,\ldots,q$. Therefore, $P(G,\alpha;k)$ in (4.1) can be written as
$P(G,\alpha;k)=\sum_{H\in\mathcal{S}_{q+1}}(-1)^{|E(H)|}k^{c(H)}.$ (4.2)
Since every element $H\in\mathcal{S}_{q+1}$ is an $\alpha$-compatible spanning subgraph of $G$ and contains no broken $\alpha$-compatible cycles, we can deduce that $H$ contains no cycles. Therefore, $H$ is a forest. This implies that for any member $H\in\mathcal{S}_{q+1}$,
$|E(H)|\leq r(G)\quad\mbox{\rm~{}and~{}}\quad c(H)=|V(G)|-|E(H)|.$
Thus, $P(G,\alpha;k)$ in (4.2) can be further simplified to the following form
$P(G,\alpha;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G,\alpha)k^{|V(G)|-i},$
which finishes the proof. ∎
Furthermore, as a direct result of Theorem 4.1, we have the next corollary.
[…]
Let $\alpha$ be an admissible assigning of $G$. Then $(-1)^{|V(G)|}P(G,\alpha;-1)$ is the number of $\alpha$-compatible spanning subgraphs of $G$ that do not contain broken $\alpha$-compatible cycles with respect to a linear ordering $\prec$.
Notably, the cycle-assigning polynomial $P(G,\alpha;k)$ agrees with the chromatic polynomial $P(G;k)$ of $G$ whenever $G$ is $\alpha$-compatible. When $G$ is $\alpha$-compatible, the $\alpha$-compatible spanning subgraph, $\alpha$-compatible cycle and broken $\alpha$-compatible cycle of $G$ correspond to the ordinary spanning subgraph, cycle and broken cycle of $G$, respectively. Hence, Theorem 4.1 deduces the following Whitney’s Broken Cycle Theorem in [30] directly.
Theorem 4.2 ([30]).
Let $\prec$ be a linear ordering on $E(G)$. Write the chromatic polynomial $P(G;k)$ as $P(G;k)=\sum_{i=0}^{r(G)}(-1)^{i}w_{i}(G)k^{|V(G)|-i}$. Then each unsigned coefficient $w_{i}(G)$ is the number of spanning subgraphs of $G$ that have exactly $i$ edges and do not contain broken cycles with respect to the linear ordering $\prec$.
5 Coefficient comparisons and further explorations
In this section, we primarily apply the Generalized Broken Cycle Theorem to explore how the coefficients of cycle-assigning polynomials vary with changes in admissible assignings and to handle certain group coloring problems in graphs. We first establish a unified order-preserving relation from admissible assignings to cycle-assigning polynomials when both are naturally ordered. Let $\alpha$ be an assigning of $G$ and $\prec$ be a linear ordering on $E(G)$. For convenience, we denote by $\mathcal{C}(G,\alpha)$ and $\mathcal{C}^{\prime}(G,\alpha)$ the sets of $\alpha$-compatible cycles and broken $\alpha$-compatible cycles of $G$ with respect to $\prec$, respectively. Additionally, let $\mathcal{C}^{\prime\prime}(G,\alpha)$ represent the set of $\alpha$-compatible spanning subgraphs of $G$ that do not contain broken $\alpha$-compatible cycles with respect to $\prec$.
Theorem 5.1 (Comparison of Coefficients).
Let $\alpha,\alpha^{\prime}$ be two admissible assignings of $G$. If $\alpha(C)\leq\alpha^{\prime}(C)$ for all cycles $C\in\mathcal{C}(G)$, then the unsigned coefficients of the cycle-assigning polynomials $P(G,\alpha;k)$ and $P(G,\alpha^{\prime};k)$ satisfy
$w_{i}(G,\alpha)\leq w_{i}(G,\alpha^{\prime})\quad\mbox{ for }\quad i=0,1,\ldots,r(G).$
In particular, $w_{i}(G)=w_{i}(G,0)\leq w_{i}(G,\alpha)$ for any admissible assigning $\alpha$ of $G$.
Proof.
By Theorem 4.1, the proof reduces to showing that every member $H$ of $\mathcal{C}^{\prime\prime}(G,\alpha)$ belongs to $\mathcal{C}^{\prime\prime}(G,\alpha^{\prime})$. Given an element $H\in\mathcal{C}^{\prime\prime}(G,\alpha)$. As $H$ is an $\alpha$-compatible spanning subgraph of $G$ containing no broken $\alpha$-compatible cycles, $H$ does not contain cycles. This implies that $H$ is $\alpha^{\prime}$-compatible. On the other hand, we have $\mathcal{C}(G,\alpha)\supseteq\mathcal{C}(G,\alpha^{\prime})$ since $\alpha(C)\leq\alpha^{\prime}(C)$ for any cycle $C$ of $G$. This derives $\mathcal{C}^{\prime}(G,\alpha)\supseteq\mathcal{C}^{\prime}(G,\alpha^{\prime})$. This further deduces that $H$ does not contain broken $\alpha^{\prime}$-compatible cycles. Then, we can conclude that $H$ is an $\alpha^{\prime}$-compatible spanning subgraph of $G$ containing no broken $\alpha^{\prime}$-compatible cycles. Namely, $\mathcal{C}^{\prime\prime}(G,\alpha)\subseteq\mathcal{C}^{\prime\prime}(G,\alpha^{\prime})$. This completes the proof. ∎
When $G$ is a loopless graph, each unsigned coefficient $w_{i}(G)$ of $P(G;k)$ is a positive integer, which is easily obtained from Theorem 4.2. Following Theorem 5.1, we directly obtain the following result.
[…]
Let $\alpha$ be an admissible assigining of $G$. Then each unsigned coefficient $w_{i}(G,\alpha)$ of the cycle-assigning polynomial $P(G,\alpha;k)$ is a positive integer, unless $\alpha(e)=0$ for some loop $e$ of $G$. Moreover, $w_{0}(G,\alpha)=1$ whenever $P(G,\alpha;k)$ is not the zero polynomial.
According to the relationship $P(G,\alpha;k)=k^{c(G)}\tau(G,\alpha;k)$, it is worth noting that the coefficients of the $\alpha$-assigning polynomial $\tau(G,\alpha;k)$ shares the same properties as the coefficients of the cycle-assigning polynomial $P(G,\alpha;k)$, as outlined in Theorem 2.4, Theorem 4.1, Section 4, Theorem 5.1 and Section 5.
It is a pleasant surprise that the Generalized Broken Cycle Theorem provides a new perspective on addressing certain group coloring problems in graphs. A well-known property in [27] says that if a graph has a nowhere-zero $A$-flow, then it admits a nowhere-zero $A^{\prime}$-flow for any Abelian group $A^{\prime}$ whose order is at least the order of $A$. This does not extend to the more general concept of group connectivity observed by Jaeger et al. [10]. For a connected planar graph $G$, Jaeger et al. [10] also showed that $G$ is $A$-colorable if and only if the dual graph $G^{*}$ is $A$-connected. Naturally, there is a corresponding problem in group colorings: for any Abelian groups $A_{1}$ and $A_{2}$ with the same order, is a $A_{1}$-colorable graph also $A_{2}$-colorable?
Most recently, Langhede and Thomassen in [21] confirmed that the property holds if $A^{\prime}$ is sufficiently large compared to $A$. This further implied that $G$ is always $A$-colorable whenever $A$ has a sufficiently large order, which can be explained by the Generalized Broken Cycle Theorem. More specifically, note the basic fact that for a loopless graph $G$, $G$ is $A$-colorable for some Abelian group $A$ if and only if $P\big{(}D(G),f;A\big{)}>0$ for any $f:E(G)\to A$. This is also equivalent to saying that $G$ is $A$-colorable if and only if $P(G,\alpha_{D,f};|A|)>0$ for any $D$-admissible assigning $\alpha_{D,f}$ related to $f:E(G)\to A$. When $G$ is a loopless graph, it follows from Theorem 4.1 that $w_{0}(G,\alpha)=1$ for any admissible assigning $\alpha$ of $G$. Thus, $P(G,\alpha;|A|)>0$ holds for all admissible assigning $\alpha$ of $G$ in this case whenever the Abelian group $A$ has a large enough order.
We conclude this section with the following corollary, which is a direct consequence of the mentioned fact.
[…]
Let $A,A^{\prime}$ be two Abelian groups of the same order, $G$ be a $A^{\prime}$-colorable graph and $D,D^{\prime}\in\mathcal{O}(G)$. If each $D$-admissible assigning $\alpha_{D,f}$ of $G$ with $f:E(G)\to A$ is also a $D^{\prime}$-admissible assigning $\alpha_{D^{\prime},f^{\prime}}$ of $G$ for some $f^{\prime}:E(G)\to A^{\prime}$, then $G$ is $A$-colorable.
Proof.
Since $G$ is $A^{\prime}$-colorable, we have $P\big{(}D^{\prime}(G),f^{\prime};A^{\prime}\big{)}>0$ for any $f^{\prime}:E(G)\to A^{\prime}$. As each $D$-admissible assigning $\alpha_{D,f}$ with $f:E(G)\to A$ is a $D^{\prime}$-admissible assigning $\alpha_{D^{\prime},f^{\prime}}$ of $G$ for some $f^{\prime}:E(G)\to A^{\prime}$, it follows that for any $f:E(G)\to A$,
$P\big{(}D(G),f;A\big{)}=P(G,\alpha;|A|)=P\big{(}D^{\prime}(G),f^{\prime};A^{\prime}\big{)}>0$
via Section 2, where $\alpha=\alpha_{D,f}=\alpha_{D^{\prime},f^{\prime}}$. This implies that $G$ is $A$-colorable. ∎
Acknowledgements
We are truly grateful to the anonymous referees for their insightful comments and valuable suggestions, which significantly improved this paper. Section 3 and Section 3 are attributed to the referees’ report and Kochol’s work [14]. The author is supported by National Natural Science Foundation of China under Grant No. 12301424.